### v1

# Distributional Scaling Laws for Emergent Capabilities

Abstract
In this paper, we explore the nature of sudden breakthroughs in language model performance at scale, which stands in contrast to smooth improvements governed by scaling laws. While advocates of “emergence” view abrupt performance gains as capabilities unlocking at specific scales, others have suggested that they are produced by thresholding effects and alleviated by continuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes, particularly when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LLM populations. These insights emphasize the role of random variation in the effect of scale on LLM capabilities.
Machine Learning, ICML
1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020) that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks (Srivastava et al., 2023).
These sudden breakthroughs fuel one of the most heated debates in modern machine learning. On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LLM to acquire specific concepts that permit out-of-distribution generalization (Wei et al., 2022). On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024). Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
We posit that a breakthrough capability is often distinguished not by direct responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics. Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds. Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a, b), even at larger scales (Madaan et al., 2024). Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks. In similar settings, existing publications consider only a few training runs and report their summary statistics. By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation. Our findings can be summarized as follows:
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or continuous for different seeds, even if all models train on the same data with the same hyperparameters. On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
  • Breakthrough scaling results from bimodal performance distributions. Using length generalization as a case study, we demonstrate that different random seeds can exhibit either highly linear or highly emergent scale behavior. We connect these differences to the bimodal distribution of this compositional skill across random seeds, a property that materializes at many parameter scales. At these scales, “emergence” is a stochastic property. We observe these bimodal distributions around breakthrough scales both in small synthetic tasks and in LLM performance on MMLU.
  • When a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of a model learning that skill may be responding continuously to scale. By modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that improvements with scale can come from changes in the probability of success or in the average performance of a successful model (Section 3). In response to the distributional nature of model performances, we propose to use Wassterstein distance to measure changes in performance distribution across model scale (Section 3.3). Using this metric, we show that the unlocking of model capacity at some small threshold in fact leads to a sudden emergence of bimodality in performance distribution from a unimodal one.
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We provide two synthetic tasks in which one has well-behaved scaling behavior and another exhibits U-shaped trends when increasing the width of the network (Section 3.4), a phenomenon also observed in LLM scaling (McKenzie et al., 2022). We note that the performance of successful runs remain monotonic during ranges of inverse scaling, confirming that U-shaped curves in performance distributions are an artifact of success probability; conventional scaling describes the performance of successful runs.
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LLMs. Our findings confirm that random variation in LLMs is also bimodal on established emergent capabilities.
(a) Fixing network depth to 6 layers.
(b) Fixing network width to 8 heads per layer.
Figure 2: Histograms of exact match accuracy when generalizing to length $40$ sequences on the reverse order addition task when independently scaling width (above) and depth (below).
2 Methodology
2.1 Synthetic Tasks
Breakthrough capabilities often require compositional reasoning (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024); we consider the compositional property of length generalization (Graves et al., 2016; Kaiser & Sutskever, 2015; Lake & Baroni, 2018; Hupkes et al., 2020). Below we outline our experimental setup, with further details in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024). To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter. We scale width by increasing the number of 64-parameter heads per layer.
Task: We consider two algorithmic tasks previously studied in Zhou et al. (2024a): counting and reverse-order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for learning to count and 200 seeds for reverse-order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. (2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Reverse-Order Addition with Index Hints: Zhou et al. (2024b) show that Transformers can generalize to 10-15 digits past their training length with two modifications made to the original addition task: generating the digits of the answer in reverse order and using index hints. Examples are given in the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where the index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work (Jelassi et al., 2024; Zhou et al., 2024b). The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for reverse-order addition.
2.2 Natural Language Tasks
We extend our investigation to LLMs by testing an emergent natural language understanding task, multiple choice question answering. We test whether the performance clusters observed in emergent synthetic tasks also manifest in LLMs. Specifically, we hypothesize that emergent scaling curves in LLMs express underlying bimodal (or multimodal) performance distributions. During training, the large and diverse pretraining corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023). This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
We conduct continued pretraining on LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024). We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
Task: We consider finetuning LLMs for the MMLU benchmark (Hendrycks et al., 2021), a multiple-choice question-answering task. Achieving strong performance on MMLU requires two key abilities: (1) natural language understanding and reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter auxiliary ability which leads to emergent trends (Srivastava et al., 2023; Hu & Frank, 2024).
Model: We use the Qwen2.5 family of base models (Yang et al., 2024). To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LLM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 (Raffel et al., 2023) with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 20 reinitializations on each data mixture ratio. We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set. We use a learning rate of 1e-5 with linear decay scheduler.
3 Synthetic Experiment Results
As noted earlier, existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the reverse order addition task from Section 2.1.
3.1 Emergence is a Sign of Bimodal Variation
First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds. Following Srivastava et al. (2023), we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales. [^1] Although the fixed seed is standard practice for reporting LLM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the reverse-order addition task, performance clusters into high and low component modes at many parameter sizes. In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities. This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression. Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence.
3.2 Sudden Jumps From Gradual Distribution Shifts
When experiments report a performance metric from only one seed or a scalar value summarizing a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
Treating the distribution as a mixture of successes and failures, we can separately analyze the probability of a successful run and the performance distribution of successful runs. Both of these properties are changing continuously and gradually when the mode increases abruptly. If we restrict our analysis to the runs achieving nontrivial 20% accuracy, we see that the probability (Figure 3 (bottom left) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Section 3.3). Even at the mode breakthrough,these underlying distributional properties are only changing gradually. We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
Figure 3: Summary statistics for models generalizing to test length 40 on the reverse-order addition task. We track overall mode (top left) and overall mean (top right). While the mode exhibits a sharp increase in accuracy at a certain model scale whether scaling width or depth, the mean evolves more continuously as a result of the bimodal nature of the random variation distribution. We also take note of the part of the distribution corresponding to successful runs by plotting the fraction of runs reaching above 20% accuracy (bottom left) and the mean of such runs (bottom right). The plots for mean include 95% confidence intervals with 1000 bootstrapped samples.
Figure 4: Scaling depth and width independently for the reverse-order addition models, we show Wasserstein-L2 distance from the performance distribution of the final, largest model. We mark the emergence of bimodality based on the last scale before multiple peaks appear in Figure 2. We mark the mode breakthrough at the point where the mode in Figure 3 (top left) sharply increases because successful length generalization becomes marginally more likely than failure. Note that the sharp decrease in W2 distance—corresponding to the appearance of bimodality at depth 3 in Figure 2—occurs before the mode breakthrough. During the apparent breakthrough in the mode, the distributions shift more gradually and continuously with increases in scale.
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 4 heads per layer.
Figure 5: Histograms for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).
3.3 Bimodality Emerges Abruptly
In the previous section, we found that bimodal random distributions cause performance breakthroughs when scaling up a single training run. We also wish to analyze how the distribution itself evolves across scales; as shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where all models are unable to length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We now show, to the contrary, that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters.
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth increases from 2 to 3 layers or when model width increases from 1 head to 2 heads. This suggests that our Wasserstein distance approach can identify the moment when a new capability is unlocked. At this moment, the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point where the mode in Figure 3 (top left) increases sharply, which always occurs. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
3.4 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Figure 5) but ultimately yields a very different scaling effect: a U-shaped curve. The summary statistics in Figure 6 reveal this peculiar phenomenon, particularly in the fixed-depth trendline. This U-shaped trend in mean accuracy is similar to that observed on certain natural language tasks (Wei et al., 2023). Furthermore, this phenomenon is not simply a result of the choice of summary statistic, as the same curve describes the evolution of the distributions as a whole when measuring their W2 distance in Figure 7.
U-shaped scaling has been observed in LLMs, but its causes are not currently well-understood. When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al. (2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 6 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 6: As in Figure 3, we capture various summary statistics for test length 60 on the count task. For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case. In the trend for fixed depth (scaling width), we see a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
Figure 7: Wasserstein-L2 distance relative to final model scale when scaling depth and width for the count task. When fixing depth and scaling width, we observe that the W2 metric has an upside-down U-shape at an intermediate range of model scales.
3.5 Is Bimodality a Mirage?
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics can make apparently emergent curves into smooth curves (Srivastava et al., 2023). We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks. Are our case studies artifacts of thresholding effects? In other words, do their bimodal distributions become unimodal under continuous metrics? To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token. If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked. Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low. Our continuous metric reports the worst mistake made by the model, i.e., the minimum probability of any correct token in the sequence.
In Figure 8, we plot a histogram for this continuous metric averaged across all samples. The distribution of this metric across random seeds is often still bimodal, evidence that bimodality is not due to thresholding alone. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 8: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples. These histograms are generated for a test length of $60$ on the count task, scaling the parameter count by increasing width with a fixed model depth (top) and by increasing depth with a fixed model width (bottom). Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
Figure 9: Violin plots of last-layer reinitialized Qwen2.5 models on the MMLU benchmark. Left Models are trained on different mixtures of the C4 news subset and the MMLU auxiliary training set. With insufficient MMLU training data, Qwen2.5-0.5B fails to follow the multiple-choice format, resulting in trivial performance. As the MMLU data ratio increases, models trained on different seeds exhibit a bimodal distribution in performance. With enough MMLU examples, all models score above the random baseline. Right: For a fixed data mix, the smaller model (Qwen2.5-0.5B) is constrained by its capacity and shows a bimodal distribution in MMLU performance. In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding. However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the full MMLU capability of the original model.
4 Natural Language Results
Having documented the bimodal variation of emergent capabilities in small synthetic settings, we turn to large language models (LLMs). We focus on the MMLU dataset, where high performance emerges after the LLM learns the multiple choice format (Srivastava et al., 2023; Hu & Frank, 2024). To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LLMs before continuing to train them.
4.1 Emergence Across Data Compositions
We first examine how data composition influences the emergence of bimodal behavior in Qwen2.5-0.5B. Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s bimodal performance distribution in ways that mirror the effect of scale.
In Figure 9 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 (Raffel et al., 2023), most models (out of 20 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format (Hu & Frank, 2024). As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline. [^2] Note that Qwen2.5-0.5B base model can already following instructions 0-shot without further finetuning, achieving 39.5% performance on MMLU. By re-intializing the last layer, we removed the base model’s MMLU ability. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full capability of the the base model. Appendix D speculates further on the effect of continued training with this pre-existing capability.
4.2 Emergence Across Model Scales
Next, we show that, as in the synthetic setting, LLM model size affects the formation of performance clusters. Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 9 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020).
5 Related Works
Random variation: Even after controlling for critical hyperparameters such as learning rate, model performance remains sensitive to stochastic aspects of the training process—specifically, random initialization and the order of training examples. Prior studies have documented consistent performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018) and observed that these differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020). Dodge et al. (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variations in out-of-sample performance. Additionally, Sellam et al. (2022) demonstrated that different BERT models independently pre-training from scratch can differ in downstream task performance and social biases. Existing work has also found model runs can cluster in out-of-distribution behavior (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation. Our work connects known these random clustering effects to the phenomenon of emergence at scale.
Length generalization and compositionality with transformers: Several studies have highlighted the challenge of length generalization in transformers (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022). Various remedies have been proposed to address length generalization failures. Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024). Others modify the dataset format, either adding scratchpad or Chain-of-Thought formats (Anil et al., 2022) or incorporating padding and index hints for specific arithmetic tasks (Jelassi et al., 2023; Zhou et al., 2024a). Regarding random variation, Zhou et al. (2024a) and Zhou et al. (2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
Emergent abilities of LLMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020). These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023). Understanding the conditions and mechanisms underlying emergence is a key area of research. Recent studies suggest that emergent abilities may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Schaeffer et al., 2024). Nonetheless, some breakthrough capabilities remain emergent. One direction to better understand the mechanisms underlying emergence is through studying language models trained on algorithmic tasks exhibiting similar behavior; for instance, Gopalani et al. (2024) show that BERT models learn the low-rank matrix completion problem with a sudden drop in the loss before interpreting components of the model before and after this transition. Snell et al. (2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. In a similar vein, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
Depth versus Width Scaling: Despite scaling laws offering a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022). For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024). Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. (2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. (2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
6 Discussion and Conclusions
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale. Previous work has already documented the variability of length generalization across random seeds (Zhou et al., 2024b, a); in general, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution performance. We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics (Juneja et al., 2023) and even in the timing of generalization breakthroughs during training (Hu et al., 2023). The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute. By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters. In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research environment increasingly concerned with compositional and breakthrough capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs. Our synthetic case study on U-shaped scaling can also enable future research into the origins of inverse scaling trends.
Future work should rigorously verify these findings at larger scales and across more diverse tasks. We especially hope future research can confirm that emergent benchmarks are bimodal when pretraining from scratch across different seeds.
Impact Statement
This work challenges conventional views on emergent capabilities in large language models by demonstrating that performance breakthroughs can be explained by multimodal distributions across training seeds rather than sharp phase transitions. By highlighting the role of random variation in scaling laws, our findings provide a new perspective on model evaluation and reproducibility, emphasizing the need for robust statistical analysis in benchmarking. This research has broad implications for interpreting model capabilities and the responsible development of AI systems.
Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. RZ, TQ, and SK acknowledge support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. RZ is supported by a Kempner Institute Graduate Research Fellowship, Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. TQ and DAM are partially supported by the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship.
Appendix A Additional Experimental Details
A.1 Synthetic Tasks
Below we provide more details when training decoder-only language models on the count and reverse order addition—with index hints— tasks. Hyperparameters are largely adapted from (Zhou et al., 2024a). We train all of our models to convergence on the train distribution.
Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
  • Model scales: As mentioned in Section 2, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
  • Hyperparameters: We use a learning rate of $31e-$ 3 with a cosine decay scheduler and weight decay 0.1. We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
Reverse Order Addition with Index Hints: For evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.
  • Model scales: For our fixed depth experiments, we fix the network depth to 6 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768\}$. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,3,4,6,8,10,12\}$ layers.
  • Hyperparameters: We use a learning rate of $41e-$ 4 with a cosine decay scheduler and weight decay 0. We set the maximum training duration to be 30000 steps, with batch size 64 and context length 512.
Appendix B Breakthroughness and Linearity
Srivastava et al. (2023) introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})},$
$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$
where $I(y)=\mathrm{sign}(\arg\max_{i}y_{i}-\arg\min_{i}y_{i})(\max_{i}y_{i}-\min_{i}y_{i})$.
Appendix C Additional Figures and Experimental Results
In Figure 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and reverse order addition (below).
Our plots in the main paper use a fixed test length 60 for count and fixed test length 40 for reverse order addition. In Figure 11 and Figure 12 we show violin plots to view the random variation distribution at each model scale across multiple test lengths for count and reverse order addition respectively.
In Section 3.3, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task. For example, Figure 13 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
In Figure 8, we saw that for the count task, the bimodal random variation distribution persists when viewing a continuous performance metric: the minimum probability averaged across all test sequences. We show in Figure 14 that a similar outcome holds for the reverse order addition task.
(a) Count.
(b) Reverse order addition.
Figure 10: Top five seeds corresponding to calculations in breakthroughness and linearity given in Appendix B for both the count task (top) and reverse order addition task (bottom).
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 11: Violin plots for the same histograms from Figure 5, i.e. for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 12: Violin plots for the same histograms from Figure 2, i.e. for EM accuracy for test length $40$ on the reverse order addition task, across 5 different scales for independently scaling depth (above) and 5 different scales for scaling width (below).
Figure 13: Histogram for EM accuracy for test length 30 (blue, i.e. in-distribution) and test length 60 (orange, i.e. out-of-distribution) on the count task when fixing depth to be one layer. While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
(a) Fixing network depth to 6 layers.
(b) Fixing network width to 8 heads per layer.
Figure 14: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples. These histograms are generated for a test length of $40$ on the reverse order addition task task, considering 5 different scaling factors for depth (top) and 3 different scaling factors for width (bottom). As we saw in Figure 8, the bimodal nature of the random variation distribution persists even when measuring probability.
Appendix D Additional LLM Results
We repeat the same data mix experiment detailed in Section 2.2 on Qwen2.5-1.5B and report results in Figure 15. In contrast to Figure 9 left, a larger model can learn to perform well on the MMLU task with fewer samples in the training data. For example, Qwen2.5-1.5B can consistently achieve non-trivial performance when the continually pretrained on a data mix that only contains 5% of MMLU auxiliary training samples. In contrast, to achieve a similar performance distribution across 20 runs, Qwen2.5-0.5B requires a data mix that contains 20% of MMLU auxiliary training samples. Interestingly, while the emergence of MMLU capability for Qwen2.5-0.5B showcase bimodal behavior, such a phenomenon is not observed in Qwen2.5-1.5B. As we increase the MMLU ratio in the training data from 0.01% to 0.5%, we instead see a more continuous transition from 0% MMLU performance to non-trivial performance. We qualitatively observe that even with a very small amount of MMLU samples, the last-layer reinitialized Qwen2.5-1.5B model can sometimes learn to output multiple-choice answers but not consistently across all the MMLU test questions. The inconsistent format following behaviors leads to a more continous transition that we observe in Figure 15.
Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup. We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model. There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments (Chen et al., 2024).
Figure 15: Violin plots for last-layer reinitialized Qwen2.5-1.5B models on the MMLU benchmark. The emergence of MMLU capability on Qwen2.5-1.5B shows a more continuous transition.

### v2

@@
 This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
-We conduct continued pretraining on LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024).
+We train LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024).
+Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences.
 We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
@@
 3 Synthetic Experiment Results
-As noted earlier, existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs.
+Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs.
 The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise.
@@
 To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth.
-In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth increases from 2 to 3 layers or when model width increases from 1 head to 2 heads.
-This suggests that our Wasserstein distance approach can identify the moment when a new capability is unlocked.
-At this moment, the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes.
+In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads.
+These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes.
 We posit that this transition marks the minimum capacity required to learn the task.
-We also mark the point where the mode in Figure 3 (top left) increases sharply, which always occurs.
+We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply.
 We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
@@
 For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case.
-In the trend for fixed depth (scaling width), we see a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
+The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
 Figure 7: Wasserstein-L2 distance relative to final model scale when scaling depth and width for the count task.
@@
 In Figure 8, we plot a histogram for this continuous metric averaged across all samples.
-The distribution of this metric across random seeds is often still bimodal, evidence that bimodality is not due to thresholding alone.
+The distribution of this metric across random seeds is often still clustered; its bimodality is not due to thresholding alone.
 We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
@@
 In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding.
-However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the full MMLU capability of the original model.
+However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the multiple choice capability of the original model.
 4 Natural Language Results
@@
 TQ and DAM are partially supported by the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship.
+Our work has been improved by invaluable discussion with Will Merrill and David Chiang.
 Appendix A Additional Experimental Details

### v3

@@
-Distributional Scaling Laws for Emergent Capabilities
+Distributional Scaling of Emergent Capabilities
 Abstract
-In this paper, we explore the nature of sudden breakthroughs in language model performance at scale, which stands in contrast to smooth improvements governed by scaling laws.
-While advocates of “emergence” view abrupt performance gains as capabilities unlocking at specific scales, others have suggested that they are produced by thresholding effects and alleviated by continuous metrics.
-We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes, particularly when performance is bimodally distributed across random seeds.
+This paper explores the nature of sudden breakthroughs in language model performance at scale, which stand in contrast to smooth improvements governed by scaling laws.
+While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics.
+We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds.
 In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends.
@@
 Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically.
-We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LLM populations.
-These insights emphasize the role of random variation in the effect of scale on LLM capabilities.
-Machine Learning, ICML
+We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations.
+These insights emphasize the role of random variation in the effect of scale on LM capabilities.
 1 Introduction
-On most benchmarks, language model (LM) performance is determined by a scaling law (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020) that varies smoothly with parameter size and overall training compute.
-There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks (Srivastava et al., 2023).
+On most benchmarks, language model (LM) performance is determined by a scaling law [12, 35, 21] that varies smoothly with parameter size and overall training compute.
+There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks [41].
 These sudden breakthroughs fuel one of the most heated debates in modern machine learning.
-On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LLM to acquire specific concepts that permit out-of-distribution generalization (Wei et al., 2022).
-On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024).
+On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LM to acquire specific concepts that permit out-of-distribution generalization [45].
+On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent [36].
 Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
-We posit that a breakthrough capability is often distinguished not by direct responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics.
+We posit that a breakthrough capability is often distinguished not by deterministic responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics.
 Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds.
-Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a, b), even at larger scales (Madaan et al., 2024).
+Although random variation may be benign when model performance is measured in-distribution [18], previous work suggests that out-of-distribution performance may vary widely across training runs [49, 51], even at larger scales [26].
 Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks.
@@
 Figure 1: Different random seeds produce different scaling trends.
-Scaling trends can be emergent or continuous for different seeds, even if all models train on the same data with the same hyperparameters.
+Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters.
 On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right).
@@
   • Breakthrough scaling results from bimodal performance distributions.
-Using length generalization as a case study, we demonstrate that different random seeds can exhibit either highly linear or highly emergent scale behavior.
-We connect these differences to the bimodal distribution of this compositional skill across random seeds, a property that materializes at many parameter scales.
+In language modeling and synthetic length generalization case studies, we demonstrate that some random seeds produce linear scale trends while others are emergent.
+This variation is caused by the bimodal distribution of each skill across random seeds, a property that materializes around breakthrough thresholds in model size.
 At these scales, “emergence” is a stochastic property.
-We observe these bimodal distributions around breakthrough scales both in small synthetic tasks and in LLM performance on MMLU.
-  • When a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of a model learning that skill may be responding continuously to scale.
-By modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that improvements with scale can come from changes in the probability of success or in the average performance of a successful model (Section 3).
-In response to the distributional nature of model performances, we propose to use Wassterstein distance to measure changes in performance distribution across model scale (Section 3.3).
-Using this metric, we show that the unlocking of model capacity at some small threshold in fact leads to a sudden emergence of bimodality in performance distribution from a unimodal one.
+We confirm these capabilities remain bimodally distributed under continuous metrics.
+  • While a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of learning that skill may be responding continuously to scale.
+Modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful model (Section 3).
+Measuring the Wassterstein distances between distributions at each scale (Section 3.3), we show that bimodality appears abruptly at some minimum capacity scale, but these sudden changes do not necessarily align with emergent trends for individual model runs.
   • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood.
-We provide two synthetic tasks in which one has well-behaved scaling behavior and another exhibits U-shaped trends when increasing the width of the network (Section 3.4), a phenomenon also observed in LLM scaling (McKenzie et al., 2022).
-We note that the performance of successful runs remain monotonic during ranges of inverse scaling, confirming that U-shaped curves in performance distributions are an artifact of success probability; conventional scaling describes the performance of successful runs.
+We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 3.4), a phenomenon also observed in LMs [28].
+We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
   • In realistic settings, multimodal performance clusters emerge based on data composition and model scale.
-In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LLMs.
-Our findings confirm that random variation in LLMs is also bimodal on established emergent capabilities.
+In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LMs.
+Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
 (a) Fixing network depth to 6 layers.
 (b) Fixing network width to 8 heads per layer.
-Figure 2: Histograms of exact match accuracy when generalizing to length $40$ sequences on the reverse order addition task when independently scaling width (above) and depth (below).
-2 Methodology
+Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length $40$ sequences when independently scaling 2(a) width and 2(b) depth.
+2 Experiments
 2.1 Synthetic Tasks
-Breakthrough capabilities often require compositional reasoning (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024); we consider the compositional property of length generalization (Graves et al., 2016; Kaiser & Sutskever, 2015; Lake & Baroni, 2018; Hupkes et al., 2020).
-Below we outline our experimental setup, with further details in Appendix A.
-Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024).
+We measure models’ length generalization [10, 20, 23, 15], one of many compositional properties that lead to emergence [41, 25, 3]. Experimental setup is described below, with details in Appendix A.
+Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) [42].
 To observe the random performance distribution at each scale, we train our models from hundreds of seeds.
@@
 Task: We consider two algorithmic tasks previously studied in Zhou et al.
-(2024a): counting and reverse-order addition.
+[49]: counting and reverse order addition.
 Performance on these task ranges widely across random seeds.
-At each scale, we train models from 250 seeds for learning to count and 200 seeds for reverse-order addition.
+At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
   • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number.
@@
 Zhou et al.
-(2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
-  • Reverse-Order Addition with Index Hints: Zhou et al.
-(2024b) show that Transformers can generalize to 10-15 digits past their training length with two modifications made to the original addition task: generating the digits of the answer in reverse order and using index hints.
-Examples are given in the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where the index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
-Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work (Jelassi et al., 2024; Zhou et al., 2024b).
-The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for reverse-order addition.
+[49] showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
+  • Reverse Order Addition with Index Hints: Zhou et al.
+[51] showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards.
+Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
+In this work, we refer to this task as the ‘addition task’ despite these modifications.
+Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work [17, 51].
+The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for addition.
+Length generalization is then tested at length 60 for count and 40 for addition.
 2.2 Natural Language Tasks
-We extend our investigation to LLMs by testing an emergent natural language understanding task, multiple choice question answering.
-We test whether the performance clusters observed in emergent synthetic tasks also manifest in LLMs.
-Specifically, we hypothesize that emergent scaling curves in LLMs express underlying bimodal (or multimodal) performance distributions.
-During training, the large and diverse pretraining corpus encourages models to acquire various capabilities.
-While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023).
-This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
-We train LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024).
+We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task [41, 40].
+(A second task, CommonSenseQA, is in Appendix E.) We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs.
+Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
 Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences.
 We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
-Task: We consider finetuning LLMs for the MMLU benchmark (Hendrycks et al., 2021), a multiple-choice question-answering task.
-Achieving strong performance on MMLU requires two key abilities: (1) natural language understanding and reasoning with domain-specific knowledge and (2) producing correct answers in the required format.
-It is this latter auxiliary ability which leads to emergent trends (Srivastava et al., 2023; Hu & Frank, 2024).
-Model: We use the Qwen2.5 family of base models (Yang et al., 2024).
-To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LLM head.
+During pretraining, a large diverse corpus encourages models to acquire various capabilities.
+While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete [29].
+This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
+Task: We consider finetuning LMs for the MMLU benchmark [11], a multiple-choice question-answering task.
+Achieving strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format.
+It is this latter ability which leads to emergent trends [41, 13].
+Model: We use the Qwen2.5 family of base models [47].
+To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head.
 We perform full-parameter continued pretraining on these reinitialized models.
-Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 (Raffel et al., 2023) with the official MMLU auxiliary training data.
+Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 [34] with the official MMLU auxiliary training data.
 The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering.
@@
 Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs.
-The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise.
+The literature suggests that these emergent behaviors are unlocked at certain model scales [45], implying that scaling curves for different model runs would generally perform similarly to within a margin of noise.
 Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
-We will introduce these concepts and our framework through a case study of the reverse order addition task from Section 2.1.
+We will introduce these concepts and our framework through a case study of the addition task from Section 2.1.
 3.1 Emergence is a Sign of Bimodal Variation
@@
 Following Srivastava et al.
-(2023), we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics.
+[41], we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics.
 As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale.
 We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales.
-[^1] Although the fixed seed is standard practice for reporting LLM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales.
+[^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales.
 As seen in Figure 1 and Appendix Figure 8, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
 This variation is explained by the bimodality of model performance distributions when varying seed.
-The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the reverse-order addition task, performance clusters into high and low component modes at many parameter sizes.
+The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the addition task, performance clusters into high and low component modes at many parameter sizes.
 In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities.
@@
 3.2 Sudden Jumps From Gradual Distribution Shifts
-When experiments report a performance metric from only one seed or a scalar value summarizing a few seeds, the outcome is likely to be close to the mode performance of the underlying model population.
+When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population.
 As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend.
@@
 We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
-Figure 3: Summary statistics for models generalizing to test length 40 on the reverse-order addition task.
-We track overall mode (top left) and overall mean (top right).
-While the mode exhibits a sharp increase in accuracy at a certain model scale whether scaling width or depth, the mean evolves more continuously as a result of the bimodal nature of the random variation distribution.
-We also take note of the part of the distribution corresponding to successful runs by plotting the fraction of runs reaching above 20% accuracy (bottom left) and the mean of such runs (bottom right).
-The plots for mean include 95% confidence intervals with 1000 bootstrapped samples.
-Figure 4: Scaling depth and width independently for the reverse-order addition models, we show Wasserstein-L2 distance from the performance distribution of the final, largest model.
-We mark the emergence of bimodality based on the last scale before multiple peaks appear in Figure 2.
-We mark the mode breakthrough at the point where the mode in Figure 3 (top left) sharply increases because successful length generalization becomes marginally more likely than failure.
-Note that the sharp decrease in W2 distance—corresponding to the appearance of bimodality at depth 3 in Figure 2—occurs before the mode breakthrough.
-During the apparent breakthrough in the mode, the distributions shift more gradually and continuously with increases in scale.
-(a) Fixing network depth to 4 layers.
-(b) Fixing network width to 4 heads per layer.
-Figure 5: Histograms for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).
+Figure 3: Summary statistics for length generalization (addition task).
+Exact match statistics for 200 models trained at length 35 and tested at length 40.
+We track overall 3 mode and 3 mean.
+Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously.
+Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs .
+Means feature 95% confidence intervals over 1000 bootstrapped samples.
+(a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
+(b) Count task. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Figure 5.
+Figure 4: Changes in random variation.
+Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently.
+We mark the emergence of bimodality at the last scale before multiple peaks appear.
+We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
 3.3 Bimodality Emerges Abruptly
-In the previous section, we found that bimodal random distributions cause performance breakthroughs when scaling up a single training run.
-We also wish to analyze how the distribution itself evolves across scales; as shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where all models are unable to length generalize.
+The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds.
+We next analyze how the distribution itself evolves across scales.
+As shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize.
 Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs.
 A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster.
-We now show, to the contrary, that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters.
+We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters.
+In Appendix Figure 10 we show that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy).
 To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth.
-In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads.
+In Figure 4(a) we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads.
 These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes.
@@
 3.4 Count: An Instance of U-shaped Scaling
-We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Figure 5) but ultimately yields a very different scaling effect: a U-shaped curve.
-The summary statistics in Figure 6 reveal this peculiar phenomenon, particularly in the fixed-depth trendline.
-This U-shaped trend in mean accuracy is similar to that observed on certain natural language tasks (Wei et al., 2023).
-Furthermore, this phenomenon is not simply a result of the choice of summary statistic, as the same curve describes the evolution of the distributions as a whole when measuring their W2 distance in Figure 7.
-U-shaped scaling has been observed in LLMs, but its causes are not currently well-understood.
-When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al.
-(2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
+We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 9) but ultimately yields a very different scaling effect: a U-shaped curve.
+Figure 5 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed.
+This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 4(b).
+U-shaped scaling has been observed in LMs, but its causes are not currently well-understood.
+When the Inverse Scaling Prize [28] solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al.
+[46] revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
 Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution.
@@
 Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
-Figure 6: As in Figure 3, we capture various summary statistics for test length 60 on the count task.
-For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case.
-The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
-Figure 7: Wasserstein-L2 distance relative to final model scale when scaling depth and width for the count task.
-When fixing depth and scaling width, we observe that the W2 metric has an upside-down U-shape at an intermediate range of model scales.
+Figure 5: Summary statistics for length generalization (count task).
+Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently.
+The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 5; however, the mean of successful runs (defined as >50% EM accuracy) 5 still exhibits continuous improvement.
 3.5 Is Bimodality a Mirage?
-Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics can make apparently emergent curves into smooth curves (Srivastava et al., 2023).
+Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior [36]; conversely, continuous metrics can make apparently emergent curves into smooth curves [41].
 We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks.
@@
 Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low.
-Our continuous metric reports the worst mistake made by the model, i.e., the minimum probability of any correct token in the sequence.
-In Figure 8, we plot a histogram for this continuous metric averaged across all samples.
-The distribution of this metric across random seeds is often still clustered; its bimodality is not due to thresholding alone.
-We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
-(a) Fixing network depth to 4 layers.
-(b) Fixing network width to 256 hidden dimension.
-Figure 8: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples.
-These histograms are generated for a test length of $60$ on the count task, scaling the parameter count by increasing width with a fixed model depth (top) and by increasing depth with a fixed model width (bottom).
-Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
-Figure 9: Violin plots of last-layer reinitialized Qwen2.5 models on the MMLU benchmark.
-Left Models are trained on different mixtures of the C4 news subset and the MMLU auxiliary training set.
-With insufficient MMLU training data, Qwen2.5-0.5B fails to follow the multiple-choice format, resulting in trivial performance.
-As the MMLU data ratio increases, models trained on different seeds exhibit a bimodal distribution in performance.
-With enough MMLU examples, all models score above the random baseline.
-Right: For a fixed data mix, the smaller model (Qwen2.5-0.5B) is constrained by its capacity and shows a bimodal distribution in MMLU performance.
-In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding.
-However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the multiple choice capability of the original model.
+We now consider a continuous version of EM accuracy that reports the worst mistake made by the model.
+This continuous error score for model $f$ on dataset $X$ is based on the per-token loss $L(f(x_{0\ldots i-1}),x_{i})$:
+$\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
+In Figure 6, we plot this continuous metric for addition models on the length generalization test dataset.
+The distribution of this metric across random seeds is often still clustered.
+We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone.
+Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
+(a) Fixing network depth to 6 layers while scaling up width.
+(b) Fixing network width to 8 heads per layer while scaling up depth.
+Figure 6: Random variation in continuous length generalization error (addition task).
+Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs.
+At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
+(a) Qwen2.5-0.5B trained on different mixes.
+(b) Qwen2.5 models trained on 10% MMLU data.
+Figure 7: Performance on MMLU.
+All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU.
+Left: With insufficient MMLU training data, each model fails to follow the multiple-choice format, resulting in trivial performance.
+As more MMLU data is included, different seeds lead to bimodally distributed performance.
+With enough MMLU data, all models score above the random baseline.
+Right: For a fixed data mix, the smaller model has bimodally distributed performance.
+In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
 4 Natural Language Results
-Having documented the bimodal variation of emergent capabilities in small synthetic settings, we turn to large language models (LLMs).
-We focus on the MMLU dataset, where high performance emerges after the LLM learns the multiple choice format (Srivastava et al., 2023; Hu & Frank, 2024).
-To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LLMs before continuing to train them.
+After connecting emergence with bimodality in small synthetic settings, we turn to LMs.
+We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format [41, 13].
+To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them.
+Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal.
+Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
 4.1 Emergence Across Data Compositions
@@
 Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s bimodal performance distribution in ways that mirror the effect of scale.
-In Figure 9 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 (Raffel et al., 2023), most models (out of 20 seeds) achieve near 0% performance on the MMLU test set.
-According to the existing literature, this failure stems from a failure to process the multiple-choice format (Hu & Frank, 2024).
+In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 [34], most models (out of 20 seeds) achieve near 0% performance on the MMLU test set.
+According to the existing literature, this failure stems from a failure to process the multiple-choice format [13].
 As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters.
@@
 Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline.
-[^2] Note that Qwen2.5-0.5B base model can already following instructions 0-shot without further finetuning, achieving 39.5% performance on MMLU.
-By re-intializing the last layer, we removed the base model’s MMLU ability.
-Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full capability of the the base model.
-Appendix D speculates further on the effect of continued training with this pre-existing capability.
 4.2 Emergence Across Model Scales
-Next, we show that, as in the synthetic setting, LLM model size affects the formation of performance clusters.
+Next, we show that, as in the synthetic setting, LM model size affects the formation of performance clusters.
 Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 7 (right).
@@
 We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales.
-This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020).
+This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models [35, 21].
+In Appendix E.1, we show similar results in the Commonsense QA [43] task.
 5 Related Works
-Random variation: Even after controlling for critical hyperparameters such as learning rate, model performance remains sensitive to stochastic aspects of the training process—specifically, random initialization and the order of training examples.
-Prior studies have documented consistent performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018) and observed that these differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020).
+Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order.
+Prior studies have documented performance differences across various stress test sets [4, 30].
+Such differences persist throughout training, not just at the final checkpoint [50].
 Dodge et al.
-(2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variations in out-of-sample performance.
+[6] compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance.
 Additionally, Sellam et al.
-(2022) demonstrated that different BERT models independently pre-training from scratch can differ in downstream task performance and social biases.
-Existing work has also found model runs can cluster in out-of-distribution behavior (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation.
-Our work connects known these random clustering effects to the phenomenon of emergence at scale.
-Length generalization and compositionality with transformers: Several studies have highlighted the challenge of length generalization in transformers (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022).
-Various remedies have been proposed to address length generalization failures.
-Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024).
-Others modify the dataset format, either adding scratchpad or Chain-of-Thought formats (Anil et al., 2022) or incorporating padding and index hints for specific arithmetic tasks (Jelassi et al., 2023; Zhou et al., 2024a).
+[38] demonstrated that LMs independently pre-trained from scratch can differ in downstream task performance and social biases.
+Existing work has also found model runs can cluster in out-of-distribution behavior [19] and in training dynamics [33, 14], hinting at multimodal variation.
+Our work connects these random clustering effects to the phenomenon of emergence at scale.
+Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on [1, 5, 9, 15, 37, 48].
+Various proposals have, consequently, aimed to improve length generalization.
+Some methods focus on alternative positional encodings [39, 32, 42, 22, 17].
+Others modify the dataset format by adding scratchpad or Chain-of-Thought [1] or by incorporating padding and index hints [16, 49].
 Regarding random variation, Zhou et al.
-(2024a) and Zhou et al.
-(2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
-Emergent abilities of LLMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020).
-These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023).
+[49] and Zhou et al.
+[51] provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
+Emergent abilities of LMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets [12, 35, 2, 21].
+These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales [45, 8, 41].
 Understanding the conditions and mechanisms underlying emergence is a key area of research.
-Recent studies suggest that emergent abilities may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Schaeffer et al., 2024).
-Nonetheless, some breakthrough capabilities remain emergent.
-One direction to better understand the mechanisms underlying emergence is through studying language models trained on algorithmic tasks exhibiting similar behavior; for instance, Gopalani et al.
-(2024) show that BERT models learn the low-rank matrix completion problem with a sudden drop in the loss before interpreting components of the model before and after this transition.
+Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale [41, 36], some breakthrough capabilities remain emergent.
 Snell et al.
-(2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus.
-In a similar vein, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
-Depth versus Width Scaling: Despite scaling laws offering a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022).
-For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024).
+[40] found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus.
+Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
+Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size [44].
+For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model [31].
 Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al.
-(2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently.
+[7] show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently.
 Levine et al.
-(2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns.
+[24] use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns.
 In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
@@
 Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale.
-Previous work has already documented the variability of length generalization across random seeds (Zhou et al., 2024b, a); in general, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution performance.
-We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics (Juneja et al., 2023) and even in the timing of generalization breakthroughs during training (Hu et al., 2023).
+Previous work has already documented the variability of length generalization across random seeds [51, 49]; in general, out-of-distribution behavior like compositional rules [27] or associative biases [38] often exhibit extreme variation compared to in-distribution performance.
+We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics [19] and even in the timing of generalization breakthroughs during training [14].
 The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution.
@@
 In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance.
-In a research environment increasingly concerned with compositional and breakthrough capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.
-Our synthetic case study on U-shaped scaling can also enable future research into the origins of inverse scaling trends.
-Future work should rigorously verify these findings at larger scales and across more diverse tasks.
-We especially hope future research can confirm that emergent benchmarks are bimodal when pretraining from scratch across different seeds.
-Impact Statement
-This work challenges conventional views on emergent capabilities in large language models by demonstrating that performance breakthroughs can be explained by multimodal distributions across training seeds rather than sharp phase transitions.
-By highlighting the role of random variation in scaling laws, our findings provide a new perspective on model evaluation and reproducibility, emphasizing the need for robust statistical analysis in benchmarking.
-This research has broad implications for interpreting model capabilities and the responsible development of AI systems.
+In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.
 Acknowledgments
@@
 Our work has been improved by invaluable discussion with Will Merrill and David Chiang.
-Appendix A Additional Experimental Details
-A.1 Synthetic Tasks
-Below we provide more details when training decoder-only language models on the count and reverse order addition—with index hints— tasks.
-Hyperparameters are largely adapted from (Zhou et al., 2024a).
+Appendix A Experimental details on synthetic tasks
+Below we provide more details when training decoder-only language models on the count and addition—with index hints— tasks.
+Hyperparameters are largely adapted from [49].
 We train all of our models to convergence on the train distribution.
+Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
 Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
@@
 Srivastava et al.
-(2023) introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale.
+[41] introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale.
 Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
@@
 where $I(y)=\mathrm{sign}(\arg\max_{i}y_{i}-\arg\min_{i}y_{i})(\max_{i}y_{i}-\min_{i}y_{i})$.
-Appendix C Additional Figures and Experimental Results
-In Figure 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and reverse order addition (below).
-Our plots in the main paper use a fixed test length 60 for count and fixed test length 40 for reverse order addition.
-In Figure 11 and Figure 12 we show violin plots to view the random variation distribution at each model scale across multiple test lengths for count and reverse order addition respectively.
+In Figure 8 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and addition (below).
+(a) Count.
+(b) Addition.
+Figure 8: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 8(a) count and 8(b) addition.
+Appendix C Additional Exact Match Accuracy Histogram Plots
+(a) Fixing network depth to 4 layers.
+(b) Fixing network width to 4 heads per layer.
+Figure 9: Random variation in length generalization (count task). Histograms of exact match accuracy on length $60$ sequences when independently scaling 9(a) width and 9(b) depth.
+Figure 10: Histograms for EM accuracy when fixing depth to be one layer (count task).
+We plot the random variation distribution at test length 30 (blue, i.e.
+in-distribution) and test length 60 (orange, i.e.
+out-of-distribution).
+While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
+Our plots in the main paper fixed test length 40 for addition (Figure 2).
+In Figure 9 we show analogous histograms for a fixed test length of 60 on the count task.
+We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
 In Section 3.3, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task.
 For example, Figure 10 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
-In Figure 8, we saw that for the count task, the bimodal random variation distribution persists when viewing a continuous performance metric: the minimum probability averaged across all test sequences.
-We show in Figure 14 that a similar outcome holds for the reverse order addition task.
-(a) Count.
-(b) Reverse order addition.
-Figure 10: Top five seeds corresponding to calculations in breakthroughness and linearity given in Appendix B for both the count task (top) and reverse order addition task (bottom).
+Appendix D Additional Continuous Metric Plots
+In Section 3.5 we have seen seen that the addition histograms are bimodal in exact match accuracy (Figure 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Figure 6).
+We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
+$\textrm{minprob}(f,X)=\frac{1}{|X|}\sum_{x\in X}\min_{i<|x|}\mathbf{P}_{f}[x_{i}|x_{0\ldots i-1}]$ (2)
+We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures.
+However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt.
+Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric.
+We plot the resulting histograms of this metric in Figures 11 and 12 for addition and count respectively.
+We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
+In Figure 13, we plot the analogous KDE plot of the loss-based error metric for the count task.
+For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition.
+However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 9 and 12 is still associated with stronger bimodality in the loss plots.
+For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions.
+For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
+(a) Fixing network depth to 6 layers while scaling up width.
+(b) Fixing network width to 8 heads per layer while scaling up depth.
+Figure 11: Histograms of the minimum probability of any token in each sequence, averaged across sequences (addition task).
+Metrics are calculated by Equation 2.
+The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Figure 6.
 (a) Fixing network depth to 4 layers.
 (b) Fixing network width to 256 hidden dimension.
-Figure 11: Violin plots for the same histograms from Figure 5, i.e.
-for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).
-(a) Fixing network depth to 4 layers.
-(b) Fixing network width to 256 hidden dimension.
-Figure 12: Violin plots for the same histograms from Figure 2, i.e.
-for EM accuracy for test length $40$ on the reverse order addition task, across 5 different scales for independently scaling depth (above) and 5 different scales for scaling width (below).
-Figure 13: Histogram for EM accuracy for test length 30 (blue, i.e.
-in-distribution) and test length 60 (orange, i.e.
-out-of-distribution) on the count task when fixing depth to be one layer.
-While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
-(a) Fixing network depth to 6 layers.
-(b) Fixing network width to 8 heads per layer.
-Figure 14: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples.
-These histograms are generated for a test length of $40$ on the reverse order addition task task, considering 5 different scaling factors for depth (top) and 3 different scaling factors for width (bottom).
-As we saw in Figure 8, the bimodal nature of the random variation distribution persists even when measuring probability.
-Appendix D Additional LLM Results
+Figure 12: Histograms of the minimum probability of any token in each sequence, averaged across sequences (count task).
+Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
+(a) Fixing network depth to 4 layers while scaling up width.
+(b) Fixing network width to 6 heads per layer while scaling up depth.
+Figure 13: Random variation in continuous length generalization error (count task).
+Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs.
+At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
+Appendix E CSQA LM Results
+E.1 Additional Task
+In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA) [43] task and we include the result in Figure 14.
+Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
+Figure 14: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.
+E.2 Impact of Model Size
 We repeat the same data mix experiment detailed in Section 2.2 on Qwen2.5-1.5B and report results in Figure 15.
@@
 We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model.
-There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments (Chen et al., 2024).
+There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments [3].
 Figure 15: Violin plots for last-layer reinitialized Qwen2.5-1.5B models on the MMLU benchmark. The emergence of MMLU capability on Qwen2.5-1.5B shows a more continuous transition.

### v4

@@
-Distributional Scaling of Emergent Capabilities
+Random Scaling of Emergent Capabilities
 Abstract
-This paper explores the nature of sudden breakthroughs in language model performance at scale, which stand in contrast to smooth improvements governed by scaling laws.
+Language models famously improve under a smooth scaling law, but some specific capabilities exhibit sudden breakthroughs in performance.
 While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics.
@@
 We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds.
-Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically.
+In a case study of inverse scaling, we show that even as the probability of a successful run declines, the average performance of a successful run increases monotonically.
 We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations.
-These insights emphasize the role of random variation in the effect of scale on LM capabilities.
+Our observations hold true even under continuous loss metrics, confirming that random variation must be considered when predicting a model’s performance from its scale.
 1 Introduction
-On most benchmarks, language model (LM) performance is determined by a scaling law [12, 35, 21] that varies smoothly with parameter size and overall training compute.
-There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks [41].
-These sudden breakthroughs fuel one of the most heated debates in modern machine learning.
-On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LM to acquire specific concepts that permit out-of-distribution generalization [45].
-On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent [36].
-Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
-We posit that a breakthrough capability is often distinguished not by deterministic responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics.
-Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds.
-Although random variation may be benign when model performance is measured in-distribution [18], previous work suggests that out-of-distribution performance may vary widely across training runs [49, 51], even at larger scales [26].
-Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks.
-In similar settings, existing publications consider only a few training runs and report their summary statistics.
-By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation.
-Our findings can be summarized as follows:
+On most benchmarks, language model (LM) performance is determined by a scaling law (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020) that varies smoothly with parameter size and overall training compute.
+There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks (Srivastava et al., 2023).
+These sudden breakthroughs fuel one of the most heated debates in modern AI.
+On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales provide the capacity to learn specific concepts (Wei et al., 2022).
+On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024).
+We will argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
+In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but emergent breakthroughs do not always reflect some fixed threshold which permits the concept to be learned.
+Instead, models may learn the concept at various scales, albeit with changing probability.
+We posit that a breakthrough capability is distinguished not by deterministic responses to scale, but by multimodal random variation.
+In other words, independent training runs cluster in their performance metrics.
+This observation is undocumented because scaling laws usually plot a single training run at each scale, rarely testing multiple seeds.
+Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a; b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
+By connecting breakthrough scaling with random variation, we challenge the narratives of both the emergence and mirage camps.
+First, our results contest the position of the mirage or “loss-to-downstream” camp, which claims that effects of scale are predictable and that continuous metrics will smooth out apparently discrete concept learning.
+We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these concepts are observable even using continuous loss metrics.
+Furthermore, we complicate the narrative of the competing emergence camp, which often treats specific model sizes as distinct in their capacity.
+While we confirm that algorithms require some minimum model capacity, that capacity may not be the enormous scale at which emergence is observed.
+Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” component ultimately dominates at larger scales.
+To extrapolate downstream metrics to large scales, we must predict the likelihood of successful emergence as well as the performance of a successful run.
+Since training numerous seeds is prohibitively expensive at large scales, we study partially reinitialized LLMs and toy models.
+Whereas prior work reports summary statistics across only a few training runs, we characterize the full multimodal performance distribution.
+We find:
 Figure 1: Different random seeds produce different scaling trends.
 Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters.
-On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right).
+On the count task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right).
 We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b).
@@
 Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
-  • Breakthrough scaling results from bimodal performance distributions.
-In language modeling and synthetic length generalization case studies, we demonstrate that some random seeds produce linear scale trends while others are emergent.
-This variation is caused by the bimodal distribution of each skill across random seeds, a property that materializes around breakthrough thresholds in model size.
-At these scales, “emergence” is a stochastic property.
-We confirm these capabilities remain bimodally distributed under continuous metrics.
-  • While a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of learning that skill may be responding continuously to scale.
-Modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful model (Section 3).
-Measuring the Wassterstein distances between distributions at each scale (Section 3.3), we show that bimodality appears abruptly at some minimum capacity scale, but these sudden changes do not necessarily align with emergent trends for individual model runs.
+  • Breakthroughs result from bimodal performance distributions.
+In multiple choice (Section 3.1) and synthetic task length generalization (Section 2.1), some random seeds produce linear scale trends while others are emergent.
+This variation is caused by the bimodal distribution of each skill across random seeds (Section 2.2), a property that materializes around breakthrough thresholds in model size.
+At these scales, emergence is a stochastic property (Section 2.3).
+  • Bimodal variation persists under continuous metrics.
+In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as accuracy rather than loss.
+For our tasks, continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs.
+We confirm these loss distributions on both synthetic (Section 2.4) and natural (Section 3.4) tasks.
+  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously.
+Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Section 2.2).
+Although bimodality can appear abruptly at a minimum capacity scale (Section 2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
   • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood.
-We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 3.4), a phenomenon also observed in LMs [28].
+We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 2.6), a phenomenon also observed in LMs (McKenzie et al., 2022).
 We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
   • In realistic settings, multimodal performance clusters emerge based on data composition and model scale.
-In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LMs.
+In Section 3, we extend our analysis to the task of multiple choice question answering by training partially reinitialized LMs.
 Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
@@
 (b) Fixing network width to 8 heads per layer.
-Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length $40$ sequences when independently scaling 2(a) width and 2(b) depth.
-2 Experiments
-2.1 Synthetic Tasks
-We measure models’ length generalization [10, 20, 23, 15], one of many compositional properties that lead to emergence [41, 25, 3]. Experimental setup is described below, with details in Appendix A.
-Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) [42].
+Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length 4040 sequences when independently scaling 2(a) width and 2(b) depth.
+2 Synthetic Length Generalization Experiments
+Usually, performance breakthroughs are measured on a single model per scale or, at most, the average of a few runs.
+The literature suggests that these emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that scaling curves for different seeds would perform similarly to within a margin of noise.
+Contrary to this belief, we demonstrate that performance is sampled from a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
+2.1 Tasks and Setup
+After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional properties that can lead to conceptual breakthroughs (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024).
+Further experimental details are in Appendix A.
+Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024).
 To observe the random performance distribution at each scale, we train our models from hundreds of seeds.
-We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter.
-We scale width by increasing the number of 64-parameter heads per layer.
+We choose model sizes by separately adjusting either the width (number of 64-parameter heads per layer) or depth hyperparameter.
 Task: We consider two algorithmic tasks previously studied in Zhou et al.
-[49]: counting and reverse order addition.
+(2024a): counting and reverse order addition.
 Performance on these task ranges widely across random seeds.
@@
 Zhou et al.
-[49] showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
-  • Reverse Order Addition with Index Hints: Zhou et al.
-[51] showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards.
+(2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
+  • Addition: Zhou et al.
+(2024b) showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards.
 Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
-In this work, we refer to this task as the ‘addition task’ despite these modifications.
-Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work [17, 51].
-The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for addition.
+We use this modified form of reverse order addition with index hints as our addition task.
+Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following prior work (Jelassi et al., 2024; Zhou et al., 2024b).
+The lengths of examples are sampled uniformly from 1 to the maximum training length (30 for count and 35 for addition).
 Length generalization is then tested at length 60 for count and 40 for addition.
-2.2 Natural Language Tasks
-We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task [41, 40].
-(A second task, CommonSenseQA, is in Appendix E.) We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs.
-Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
-Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences.
-We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
-During pretraining, a large diverse corpus encourages models to acquire various capabilities.
-While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete [29].
-This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
-Task: We consider finetuning LMs for the MMLU benchmark [11], a multiple-choice question-answering task.
-Achieving strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format.
-It is this latter ability which leads to emergent trends [41, 13].
-Model: We use the Qwen2.5 family of base models [47].
-To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head.
-We perform full-parameter continued pretraining on these reinitialized models.
-Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 [34] with the official MMLU auxiliary training data.
-The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering.
-We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
-Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 20 reinitializations on each data mixture ratio.
-We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set.
-We use a learning rate of 1e-5 with linear decay scheduler.
-3 Synthetic Experiment Results
-Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs.
-The literature suggests that these emergent behaviors are unlocked at certain model scales [45], implying that scaling curves for different model runs would generally perform similarly to within a margin of noise.
-Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
-We will introduce these concepts and our framework through a case study of the addition task from Section 2.1.
 2.2 Emergence is a Sign of Bimodal Variation
-First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds.
+What do the scaling curves for length generalization tasks look like when we generate them in conventional ways?
 Following Srivastava et al.
-[41], we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics.
+(2023), we take the vector of model performances for the count task at test length 60 across different scales with a fixed initialization and shuffle seed.
+[^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales.
+We calculate the task’s breakthroughness and linearity metrics on each seed.
 As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale.
-We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales.
-[^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales.
+We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1.
 As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
 This variation is explained by the bimodality of model performance distributions when varying seed.
-The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the addition task, performance clusters into high and low component modes at many parameter sizes.
-In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities.
-This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression.
-Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster.
+Figure 2 illustrates that, for a population of models independently trained on the addition task, length generalization clusters into high and low component modes at many parameter sizes.
+This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others follow a more linear progression.
+When bimodally distributed runs cluster into very high and very low performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster.
 These differences ultimately lead to high variability in the timing and degree of emergence.
+Furthermore, these differences can be a source of oscillating scale curves, as Srivastava et al.
+(2023) documented in their lowest-linearity tasks.
+Figure 3: Summary statistics for length generalization (addition task).
+Exact match statistics for 200 models trained at length 35 and tested at length 40.
+We track overall 3 mode and 3 mean.
+Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously.
+Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs.
+Means feature 95% confidence intervals over 1000 bootstrapped samples.
 2.3 Sudden Jumps From Gradual Distribution Shifts
@@
 As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend.
-However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves.
+However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although the underlying bimodality also produces discontinuity in many single-seed scaling curves.
 Underlying this discontinuous performance jump are continuous changes in other distributional statistics.
@@
 We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
-Figure 3: Summary statistics for length generalization (addition task).
-Exact match statistics for 200 models trained at length 35 and tested at length 40.
-We track overall 3 mode and 3 mean.
-Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously.
-Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs .
-Means feature 95% confidence intervals over 1000 bootstrapped samples.
+2.4 Is Bimodality a Mirage?
+(a) Fixing network depth to 6 layers while scaling up width.
+(b) Fixing network width to 8 heads per layer while scaling up depth.
+Figure 4: Random variation in continuous length generalization error (addition task).
+Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs.
+At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
+Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023).
+We must be particularly cautious about claiming emergence when requiring outputs to exactly match a target string, as we have so far.
+Are our case studies artifacts of thresholding effects?
+Do our bimodal distributions become unimodal under continuous metrics?
+To avoid thresholding artifacts, we use a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token.
+Each token is individually computed in the correct output context, so the entire sequence represents a fixed set of continuous per-token loss functions.
+Because the maximum of a fixed set of Lipschitz-continuous functions is always Lipschitz-continuous, this error score is guaranteed to be continuous as long as per-token loss is continuous.
+For model $f$ on dataset $X$, the continuous error score based on per-token loss $L(f(x_{0\ldots i-1}),x_{i})$ is:
+$\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
+In Figure 4, we plot this continuous metric for addition models on the length generalization test dataset.
+The distribution of this metric across random seeds is often still clustered.
+We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone.
+Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
+2.5 Bimodality Emerges Abruptly
 (a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
@@
 We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
-3.3 Bimodality Emerges Abruptly
 The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds.
@@
 2.6 Count: An Instance of U-shaped Scaling
-We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 9) but ultimately yields a very different scaling effect: a U-shaped curve.
+We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 11) but yields a very different scaling effect: a U-shaped curve.
 Figure 6 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed.
@@
 U-shaped scaling has been observed in LMs, but its causes are not currently well-understood.
-When the Inverse Scaling Prize [28] solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al.
-[46] revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
+When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al.
+(2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
 Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution.
@@
 Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently.
-The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 5; however, the mean of successful runs (defined as >50% EM accuracy) 5 still exhibits continuous improvement.
-3.5 Is Bimodality a Mirage?
-Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior [36]; conversely, continuous metrics can make apparently emergent curves into smooth curves [41].
-We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks.
-Are our case studies artifacts of thresholding effects?
-In other words, do their bimodal distributions become unimodal under continuous metrics?
-To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token.
-If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked.
-Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low.
-We now consider a continuous version of EM accuracy that reports the worst mistake made by the model.
-This continuous error score for model $f$ on dataset $X$ is based on the per-token loss $L(f(x_{0\ldots i-1}),x_{i})$:
-$\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
-In Figure 6, we plot this continuous metric for addition models on the length generalization test dataset.
-The distribution of this metric across random seeds is often still clustered.
-We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone.
-Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
-(a) Fixing network depth to 6 layers while scaling up width.
-(b) Fixing network width to 8 heads per layer while scaling up depth.
-Figure 6: Random variation in continuous length generalization error (addition task).
-Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs.
-At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
+The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 6; however, the mean of successful runs (defined as >50% EM accuracy) 6 still improves continuously.
 (a) Qwen2.5-0.5B trained on different mixes.
@@
 In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
-4 Natural Language Results
+3 LLM Experiments
 After connecting emergence with bimodality in small synthetic settings, we turn to LMs.
-We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format [41, 13].
+We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format (Srivastava et al., 2023; Hu and Frank, 2024).
 To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them.
@@
 Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
+3.1 Data and Setup
+We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task (Srivastava et al., 2023; Snell et al., 2024).
+We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs.
+Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
+Our procedure can be regarded as continued pretraining: we mimic the model’s original pretraining objective with a dataset of English language sequences.
+We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
+During pretraining, a large diverse corpus encourages models to acquire various capabilities.
+While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023).
+This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
+Task: We test LMs on MMLU Hendrycks et al.
+(2021), a multiple-choice question-answering benchmark.
+Strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format.
+It is this latter ability which leads to emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
+Model: We use the Qwen2.5 family of base models Yang et al.
+(2024).
+To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head.
+We perform full-parameter continued pretraining on these reinitialized models.
+Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 Raffel et al.
+(2023) with the official MMLU auxiliary training data.
+The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering.
+We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
+Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 80 reinitializations on each data mixture ratio.
+We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set.
+We use a learning rate of 1e-5 with linear decay scheduler.
+We focus our analysis on experiment results from Qwen2.5-0.5B.
+In Appendix E, we include results from Qwen2.5-1.5B (Appendix 17) as well as a second task CommonSenseQA (Talmor et al., 2019) (Appendix E.2).
 3.2 Emergence Across Data Compositions
-We first examine how data composition influences the emergence of bimodal behavior in Qwen2.5-0.5B.
-Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s bimodal performance distribution in ways that mirror the effect of scale.
-In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 [34], most models (out of 20 seeds) achieve near 0% performance on the MMLU test set.
-According to the existing literature, this failure stems from a failure to process the multiple-choice format [13].
+We first examine how data composition influences the bimodal capabilities of Qwen2.5-0.5B.
+Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s performance distribution in ways that mirror the effect of scale.
+In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 Raffel et al.
+(2023), most models (out of 80 seeds) achieve near 0% performance on the MMLU test set.
+According to the existing literature, this failure stems from a failure to process the multiple-choice format Hu and Frank (2024).
 As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters.
@@
 This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer.
-Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline.
+Finally, with sufficient data ($20\%$ MMLU), models across all 80 seeds consistently outperform the 25% random baseline.
 3.3 Emergence Across Model Scales
-Next, we show that, as in the synthetic setting, LM model size affects the formation of performance clusters.
-Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 7 (right).
+Like in the synthetic setting, LM model size affects the clustered performance distribution.
+Using a training data mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 80 seeds and report MMLU test accuracy in Figure 7 (right).
 While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline.
-We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales.
-This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models [35, 21].
-In Appendix E.1, we show similar results in the Commonsense QA [43] task.
-5 Related Works
+In fact, larger models reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales.
+This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020).
+Appendix E.2 shows similar results on the Commonsense QA (Talmor et al., 2019) task.
+Figure 8: NLL loss distributions on the MMLU test set.
+Each distribution is provided as a histogram and as a smoothed KDE.
+Even with a continuous metric (NLL), we still observe multi-modal distribution across random seeds.
+Figure 9: KDE approximation of LM training loss on 10% MMLU mixture.
+Approximation of the distribution across 80 random seeds is unimodal, smooth, and nearly symmetrical with median close to mean.
+By contrast, Figure 8 depicts MMLU loss for LMs trained on the same mix as irregularly clustered with multiple modes.
+3.4 Is MMLU emergence a mirage?
+Is multiple choice performance only bimodal because accuracy is thresholded?
+As in our synthetic setting (Section 2.4), we must address the effect of the discontinuous metric.
+In Figure 8, we find that continuous negative log likelihood (NLL) loss metrics remain highly multimodal.
+The 10% MMLU ratio leads to the most even split between successes and failure in accuracy, and accordingly to clear bimodality in NLL loss, mirroring the link we observed on synthetic tasks.
+Notably, training data with a 7.5% MMLU ratio produces a performance distribution with a small “success” cluster peaked around NLL=3.4.
+The 10% MMLU ratio produces a performance distribution with a larger success cluster, also peaking around NLL=3.4.
+Therefore, the continuous metrics remain roughly bimodal—as is especially clear in the 10% mix setting, which is most evenly split between success and failure.
+While recent arguments against emergence (Schaeffer et al., 2024) claim that discontinuities disappear with continuous loss metrics, Figure 8 reveals the opposite effect in random variation—NLL loss reveals clusters that are concealed by accuracy metrics, specifically among low-performance models like those trained on the 5% MMLU mix.
+(To easily compare these metrics, see Figure 13.)
+Is this random clustering particular to the emergent task?
+As an alternative hypothesis, we might expect all tasks—including training loss—to have similarly multimodal NLL distributions, when approximated with our KDE settings.
+To the contrary, the pretraining loss produces a smooth, unimodal, and almost symmetric KDE distribution (Figure 9).
+We therefore conclude that multimodality is linked to emergent tasks—not merely a downstream effect of in-distribution training noise.
+4 Discussion
+Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities.
+The mode of our performance distributions sharply improves at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution.
+Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 2.5); the transition from the initial unimodal distribution to a bimodal one is sudden.
 Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order.
-Prior studies have documented performance differences across various stress test sets [4, 30].
-Such differences persist throughout training, not just at the final checkpoint [50].
+Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018).
+Most relevant to us, the high variability of length generalization is well-documented (Zhou et al., 2024b; a).
+More generally, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution loss.
+Such differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020).
 Dodge et al.
-[6] compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance.
-Additionally, Sellam et al.
-[38] demonstrated that LMs independently pre-trained from scratch can differ in downstream task performance and social biases.
-Existing work has also found model runs can cluster in out-of-distribution behavior [19] and in training dynamics [33, 14], hinting at multimodal variation.
+(2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance.
+Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation.
 Our work connects these random clustering effects to the phenomenon of emergence at scale.
-Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on [1, 5, 9, 15, 37, 48].
+Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022).
 Various proposals have, consequently, aimed to improve length generalization.
-Some methods focus on alternative positional encodings [39, 32, 42, 22, 17].
-Others modify the dataset format by adding scratchpad or Chain-of-Thought [1] or by incorporating padding and index hints [16, 49].
+Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024).
+Others modify the dataset format by adding scratchpad or Chain-of-Thought (Anil et al., 2022) or by incorporating padding and index hints (Jelassi et al., 2023; Zhou et al., 2024a).
 Regarding random variation, Zhou et al.
-[49] and Zhou et al.
-[51] provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
-Emergent abilities of LMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets [12, 35, 2, 21].
-These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales [45, 8, 41].
+(2024a) and Zhou et al.
+(2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
+Emergent abilities of LMs: In LLMs, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020).
+These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023).
 Understanding the conditions and mechanisms underlying emergence is a key area of research.
-Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale [41, 36], some breakthrough capabilities remain emergent.
+Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Srivastava et al., 2023; Schaeffer et al., 2024), some breakthrough capabilities remain emergent.
 Snell et al.
-[40] found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus.
+(2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus.
 Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
-Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size [44].
-For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model [31].
+Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022).
+For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024).
 Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al.
-[7] show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently.
+(2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently.
 Levine et al.
-[24] use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns.
+(2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns.
 In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
-6 Discussion and Conclusions
-Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale.
-Previous work has already documented the variability of length generalization across random seeds [51, 49]; in general, out-of-distribution behavior like compositional rules [27] or associative biases [38] often exhibit extreme variation compared to in-distribution performance.
-We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics [19] and even in the timing of generalization breakthroughs during training [14].
-The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution.
-Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
-In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute.
-By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters.
-In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance.
-In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.
+In one surprising result, we document a regime in which increasing width damages model performance but increasing depth improves model performance.
+In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
 Acknowledgments
@@
 Our work has been improved by invaluable discussion with Will Merrill and David Chiang.
+Reproducibility Statement
+Upon publication, we will release the code for both synthetic and LLM experiments, including code, and data.
+Detailed descriptions of model, data and training are provided in Appendix A, and Section 3.
 Appendix A Experimental details on synthetic tasks
 Below we provide more details when training decoder-only language models on the count and addition—with index hints—tasks.
-Hyperparameters are largely adapted from [49].
+Hyperparameters are largely adapted from [Zhou et al., 2024a].
 We train all of our models to convergence on the train distribution.
@@
 Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
-  • Model scales: As mentioned in Section 2, we scale up our models by fixing width and scaling depth and fixing depth and scaling width.
+  • Model scales: As mentioned in Section 2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width.
 The precise parameters for each variation are as follows.
@@
 For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
-  • Hyperparameters: We use a learning rate of $31e-$ 3 with a cosine decay scheduler and weight decay 0.1.
+  • Hyperparameters: We use a learning rate of 1e31e-3 with a cosine decay scheduler and weight decay 0.1.
 We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
@@
 For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,3,4,6,8,10,12\}$ layers.
-  • Hyperparameters: We use a learning rate of $41e-$ 4 with a cosine decay scheduler and weight decay 0.
+  • Hyperparameters: We use a learning rate of 1e41e-4 with a cosine decay scheduler and weight decay 0.
 We set the maximum training duration to be 30000 steps, with batch size 64 and context length 512.
@@
 (b) Addition.
-Figure 8: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 8(a) count and 8(b) addition.
+Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 10(a) count and 10(b) addition.
 Appendix C Additional Exact Match Accuracy Histogram Plots
@@
 (b) Fixing network width to 4 heads per layer.
-Figure 9: Random variation in length generalization (count task). Histograms of exact match accuracy on length $60$ sequences when independently scaling 9(a) width and 9(b) depth.
+Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling 11(a) width and 11(b) depth.
 Figure 12: Histograms for EM accuracy when fixing depth to be one layer (count task).
@@
 Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric.
-We plot the resulting histograms of this metric in Figures 11 and 12 for addition and count respectively.
+We plot the resulting histograms of this metric in Figures 14 and 15 for addition and count respectively.
 We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
@@
 For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition.
-However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 9 and 12 is still associated with stronger bimodality in the loss plots.
+However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 11 and 15 is still associated with stronger bimodality in the loss plots.
 For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions.
 For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
+Finally, Figure 13 provides a copy of Figure 8 with both accuracy and NLL loss comparable for easy contrast.
+We note that high-performance clusters are distinguishable in both accuracy and loss, but the thresholded accuracy conceals low-performance clusters which are visible in the continuous distribution.
+Figure 13: KDE approximations of performance variation on the MMLU test set.
+Provided for contrasting accuracy and continuous NLL loss.
+Top: Approximate distribution of accuracy metrics.
+Bottom: Approximate distribution of KDE metrics, copied from Figure 8.
 (a) Fixing network depth to 6 layers while scaling up width.
@@
 At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
-Appendix E CSQA LM Results
-E.1 Additional Task
-In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA) [43] task and we include the result in Figure 14.
+Appendix E Additional Language Model Results
+E.1 MMLU Results on Qwen2.5-1.5B
+We repeat the same MMLU experiment (Section 3.1) on Qwen2.5-1.5B and we include results in Figure 17.
+Due to computation limit, for each data mixes, we train on 20 random seeds.
+Similar to Figure 13.
+models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
+Figure 17: Accuracy and NLL loss for Qwen2.5-1.5B MMLU experiments.
+We train last-layer reinitialized Qwen2.5-1.5B on data mixes with different ratios of the target task (MMLU).
+Top: Distribution of accuracy metric.
+Bottom: Distribution of NLL loss metric.
+E.2 CSQA LM Results
+In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA, Talmor et al.
+[2019]) task and we include the result in Figure 18.
 Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
 Figure 18: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.
-E.2 Impact of Model Size
-We repeat the same data mix experiment detailed in Section 2.2 on Qwen2.5-1.5B and report results in Figure 15.
-In contrast to Figure 7 left, a larger model can learn to perform well on the MMLU task with fewer samples in the training data.
-For example, Qwen2.5-1.5B can consistently achieve non-trivial performance when the continually pretrained on a data mix that only contains 5% of MMLU auxiliary training samples.
-In contrast, to achieve a similar performance distribution across 20 runs, Qwen2.5-0.5B requires a data mix that contains 20% of MMLU auxiliary training samples.
-Interestingly, while the emergence of MMLU capability for Qwen2.5-0.5B showcase bimodal behavior, such a phenomenon is not observed in Qwen2.5-1.5B.
-As we increase the MMLU ratio in the training data from 0.01% to 0.5%, we instead see a more continuous transition from 0% MMLU performance to non-trivial performance.
-We qualitatively observe that even with a very small amount of MMLU samples, the last-layer reinitialized Qwen2.5-1.5B model can sometimes learn to output multiple-choice answers but not consistently across all the MMLU test questions.
-The inconsistent format following behaviors leads to a more continous transition that we observe in Figure 15.
-Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup.
-We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model.
-There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments [3].
-Figure 15: Violin plots for last-layer reinitialized Qwen2.5-1.5B models on the MMLU benchmark. The emergence of MMLU capability on Qwen2.5-1.5B shows a more continuous transition.

### v5

@@
 Language models famously improve under a smooth scaling law, but some specific capabilities exhibit sudden breakthroughs in performance.
-While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics.
+Advocates of “emergence” view these capabilities as unlocked at a specific scale, but others attribute breakthroughs to superficial metric thresholding effects.
 We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds.
-In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends.
+we show that different random seeds can produce either smooth or emergent scaling trends in synthetic length generalization tasks, multiple choice question answering, and grammatical generalization.
 We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds.
-In a case study of inverse scaling, we show that even as the probability of a successful run declines, the average performance of a successful run increases monotonically.
-We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations.
+These distributions may become abruptly bimodal at a capacity threshold—but this threshold appears at scales well before most seeds achieve breakthrough.
 Our observations hold true even under continuous loss metrics, confirming that random variation must be considered when predicting a model’s performance from its scale.
+Machine Learning, ICML
 1 Introduction
@@
 These sudden breakthroughs fuel one of the most heated debates in modern AI.
-On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales provide the capacity to learn specific concepts (Wei et al., 2022).
-On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024).
-We will argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
-In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but emergent breakthroughs do not always reflect some fixed threshold which permits the concept to be learned.
+On one side, advocates of emergence claim that performance abruptly improves when a particular scale provides the capacity to learn specific concepts (Wei et al., 2022).
+On the other side, skeptics argue that these sudden improvements are a mirage (Schaeffer et al., 2024) driven by thresholding effects.
+These threshold artifacts are alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent.
+We argue that discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
+In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but breakthroughs do not always reflect a fixed threshold at which a concept is learnable.
 Instead, models may learn the concept at various scales, albeit with changing probability.
+Figure 1: Different random seeds produce different scaling trends.
+Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters.
+On the count task (see Appx.
+C), we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right).
+We mark parameter counts immediately before and after seed 93’s emergence, respectively, as (a) and (b).
+Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205.
+Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
 We posit that a breakthrough capability is distinguished not by deterministic responses to scale, but by multimodal random variation.
@@
 This observation is undocumented because scaling laws usually plot a single training run at each scale, rarely testing multiple seeds.
-Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a; b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
+Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a, b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
 By connecting breakthrough scaling with random variation, we challenge the narratives of both the emergence and mirage camps.
-First, our results contest the position of the mirage or “loss-to-downstream” camp, which claims that effects of scale are predictable and that continuous metrics will smooth out apparently discrete concept learning.
-We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these concepts are observable even using continuous loss metrics.
-Furthermore, we complicate the narrative of the competing emergence camp, which often treats specific model sizes as distinct in their capacity.
+First, our results contest the position of the mirage or “loss-to-downstream” camp: that effects of scale are predictable and that continuous metrics will smooth out apparent breakthroughs.
+We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these clusters are observable even using continuous loss metrics.
+Furthermore, we complicate the rival narrative of the emergence camp, which often treats specific model sizes as distinct in their capacity.
 While we confirm that algorithms require some minimum model capacity, that capacity may not be the enormous scale at which emergence is observed.
-Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” component ultimately dominates at larger scales.
+Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” mode ultimately dominates at larger scales.
 To extrapolate downstream metrics to large scales, we must predict the likelihood of successful emergence as well as the performance of a successful run.
@@
 We find:
-Figure 1: Different random seeds produce different scaling trends.
-Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters.
-On the count task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right).
-We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b).
-Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205.
-Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
-  • Breakthroughs result from bimodal performance distributions.
-In multiple choice (Section 3.1) and synthetic task length generalization (Section 2.1), some random seeds produce linear scale trends while others are emergent.
-This variation is caused by the bimodal distribution of each skill across random seeds (Section 2.2), a property that materializes around breakthrough thresholds in model size.
-At these scales, emergence is a stochastic property (Section 2.3).
-  • Bimodal variation persists under continuous metrics.
-In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as accuracy rather than loss.
-For our tasks, continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs.
-We confirm these loss distributions on both synthetic (Section 2.4) and natural (Section 3.4) tasks.
-  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously.
-Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Section 2.2).
-Although bimodality can appear abruptly at a minimum capacity scale (Section 2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
-  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood.
-We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 2.6), a phenomenon also observed in LMs (McKenzie et al., 2022).
-We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
-  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale.
-In Section 3, we extend our analysis to the task of multiple choice question answering by training partially reinitialized LMs.
-Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
 (a) Fixing network depth to 6 layers.
@@
 Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length 4040 sequences when independently scaling 2(a) width and 2(b) depth.
-2 Synthetic Length Generalization Experiments
-Usually, performance breakthroughs are measured on a single model per scale or, at most, the average of a few runs.
-The literature suggests that these emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that scaling curves for different seeds would perform similarly to within a margin of noise.
+  • Breakthroughs result from bimodal performance distributions.
+On synthetic algorithmic tasks (Sec.
+2), simple language modeling tasks (Sec.
+4) and multiple choice question answering (Sec.
+3) , some random seeds produce linear scale trends while others are emergent.
+This variation is caused by the bimodal distribution of each skill across seeds (Sec.
+2.2), a property that materializes around breakthrough thresholds in model size.
+At these scales, emergence is a stochastic property (Sec.
+2.3).
+  • Bimodal variation persists under continuous metrics.
+In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as exact match accuracy rather than loss.
+We find continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs.
+We confirm these results on both synthetic (Sec.
+2.4, Sec.
+4.2) and natural (Sec.
+3.3) tasks.
+  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously.
+Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Sec.
+2.2).
+Although bimodality can appear abruptly at a minimum capacity scale (Sec.
+2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
+  • Bimodality can produce nonmonotonic scaling trends.
+We discuss a synthetic task with ranges of inverse scaling in mean accuracy (Appx.
+C.2).
+The average of successful runs remains monotonic in these ranges, suggesting that U-shaped curves in overall mean are an artifact of success probability.
+2 Synthetic Length Generalization Tasks
+Usually, scaling trends are measured on a single model per scale or, at most, the average of a few runs.
+The literature suggests that emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that different training runs would perform similarly at each scale.
 Contrary to this belief, we demonstrate that performance is sampled from a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
 2.1 Tasks and Setup
-After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional properties that can lead to conceptual breakthroughs (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024).
-Further experimental details are in Appendix A.
+After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional skills that can lead to conceptual breakthroughs (Srivastava et al., 2023; löwe2024abruptspontaneousstrategyswitches; Chen et al., 2024).
+Experimental details are in Appx.
+A.
 Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024).
@@
 Task: We consider two algorithmic tasks previously studied in Zhou et al.
-(2024a): counting and reverse order addition.
-Performance on these task ranges widely across random seeds.
-At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
-  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number.
-Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training.
+(2024a): counting and addition.
+An analysis of the count task is provided in Appx.
+C, including discussion (Appx.
+C.2) of how its average accuracy responds non-monotonically to scale.
+In the main body of this paper, we will focus on the addition task.
 Zhou et al.
-(2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
-  • Addition: Zhou et al.
 (2024b) showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards.
-Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
-We use this modified form of reverse order addition with index hints as our addition task.
+We use this modified reverse-order addition task.
 Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following prior work (Jelassi et al., 2024; Zhou et al., 2024b).
@@
 Following Srivastava et al.
-(2023), we take the vector of model performances for the count task at test length 60 across different scales with a fixed initialization and shuffle seed.
-[^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales.
-We calculate the task’s breakthroughness and linearity metrics on each seed.
-As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale.
-We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1.
-As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
+(2023), we calculate the breakthroughness and linearity of scaling curves where each run shares a fixed initialization and shuffle seed.
+[^1] Although it is standard practice to fix the random seed when comparing LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation between different model sizes.
+As defined in Appx.
+B, breakthroughness measures emergence, whereas linearity measures a smooth response to scale.
+We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Fig.
+1 (with other extremes shown in Appx.
+Fig.
+10).
+Because breakthroughs vary across seeds, we can easily find fixed seeds that lead to either emergent or smooth scaling.
 This variation is explained by the bimodality of model performance distributions when varying seed.
-Figure 2 illustrates that, for a population of models independently trained on the addition task, length generalization clusters into high and low component modes at many parameter sizes.
-This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others follow a more linear progression.
-When bimodally distributed runs cluster into very high and very low performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster.
+Fig.
+2 illustrates that, for a population of models independently trained on the addition task, length generalization ability clusters into high and low component modes at many parameter sizes.
+This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others generalize poorly.
+When bimodally distributed runs cluster into distinct high- and low-performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster.
 These differences ultimately lead to high variability in the timing and degree of emergence.
-Furthermore, these differences can be a source of oscillating scale curves, as Srivastava et al.
-(2023) documented in their lowest-linearity tasks.
-Figure 3: Summary statistics for length generalization (addition task).
-Exact match statistics for 200 models trained at length 35 and tested at length 40.
-We track overall 3 mode and 3 mean.
-Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously.
-Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs.
-Means feature 95% confidence intervals over 1000 bootstrapped samples.
+Furthermore, these differences cause oscillating scale curves like those seen in low-linearity tasks (Srivastava et al., 2023, ref.
+Fig.
+7c).
+Figure 3: Summary statistics for length generalization accuracy on the reverse-order addition task.
+We track overall mode (top left) and overall mean (top right).
+While the mode exhibits a sharp increase in accuracy at a certain scale of width or depth, the mean evolves more continuously as a result of the bimodal nature of the random variation distribution.
+We also study successful runs, plotting the fraction of runs reaching above 20% accuracy (bottom left) and the mean of such runs (bottom right).
+The plots for mean include 95% confidence intervals with 1000 bootstrapped samples.
 2.3 Sudden Jumps From Gradual Distribution Shifts
 When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population.
-As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend.
-However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although the underlying bimodality also produces discontinuity in many single-seed scaling curves.
+In Fig.
+3 (top left), the mode performance shows a massive spike, mirroring an emergent benchmark’s scaling trend.
+However, we claim that this discontinuous improvement in mode—as well as the discontinuities seen in single-seed emergent scaling curves—is an artifact of bimodality.
 Underlying this discontinuous performance jump are continuous changes in other distributional statistics.
-In Figure 3 (top right), the mean exhibits a smoother trend in accuracy.
-Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
+In Fig.
+3 (top right), the mean exhibits a smoother trend in accuracy.
+Mode and mean differ because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
 Treating the distribution as a mixture of successes and failures, we can separately analyze the probability of a successful run and the performance distribution of successful runs.
 Both of these properties are changing continuously and gradually when the mode increases abruptly.
-If we restrict our analysis to the runs achieving nontrivial 20% accuracy, we see that the probability (Figure 3 (bottom left) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Section 2.5).
+If we restrict our analysis to the runs achieving at least nontrivial 20% accuracy, we see that the probability (Fig.
+3 (bottom left)) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Sec.
+2.5).
 Even at the mode breakthrough, these underlying distributional properties are only changing gradually.
-We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
+We conclude that when tasks exhibit bimodal distributions across random seeds, gradual statistical improvements cause seemingly abrupt improvements with scale.
 2.4 Is Bimodality a Mirage?
-(a) Fixing network depth to 6 layers while scaling up width.
-(b) Fixing network width to 8 heads per layer while scaling up depth.
-Figure 4: Random variation in continuous length generalization error (addition task).
-Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs.
-At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
-Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023).
+Metrics with hard thresholds can artificially induce breakthroughs (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023).
 We must be particularly cautious about claiming emergence when requiring outputs to exactly match a target string, as we have so far.
@@
 Do our bimodal distributions become unimodal under continuous metrics?
-To avoid thresholding artifacts, we use a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token.
+To avoid thresholding artifacts, consider a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token.
 Each token is individually computed in the correct output context, so the entire sequence represents a fixed set of continuous per-token loss functions.
@@
 $\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
-In Figure 4, we plot this continuous metric for addition models on the length generalization test dataset.
-The distribution of this metric across random seeds is often still clustered.
-We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone.
-Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
+In Fig.
+4 (a)(b), we plot this continuous metric for addition models on the length generalization dataset.
+The distribution of this metric across random seeds is still clustered.
+We therefore confirm that emergent capabilities exhibit bimodal distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone.
+Further findings from these experiments are in Appx.
+D.
 2.5 Bimodality Emerges Abruptly
-(a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
-(b) Count task. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Figure 6.
-Figure 5: Changes in random variation.
-Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently.
-We mark the emergence of bimodality at the last scale before multiple peaks appear.
-We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
+(a) Fixing network depth to 6 layers while scaling up width.
+(b) Fixing network width to 8 heads per layer while scaling up depth.
+(c) Changes in random variation.
+Figure 4: Analysis of performance distributions (addition task).
+Left (a, b): Kernel Density Estimation (KDE) of our loss-based error metric (Eq.
+1).
+At scales where accuracy is bimodal, the distribution remains bimodal even with a continuous metric.
+Right (c): Wasserstein-L2 distance of each scale’s performance relative to the largest scale.
+We mark the “emergence of bimodality” and the “mode breakthrough.” During breakthrough, distributions shift gradually.
 The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds.
-We next analyze how the distribution itself evolves across scales.
-As shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize.
-Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs.
+We next show that the distribution itself can change suddenly at certain scales.
+As shown in Fig.
+2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize.
+Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful (i.e., length-generalizing) runs.
 A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster.
 We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters.
-In Appendix Figure 12 we show that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy).
-To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth.
-In Figure 5(a) we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads.
-These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes.
+In Appx.
+Fig.
+12, we confirm that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy), illustrating that the length-general function is suddenly learnable at sufficient depth.
+To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each scale’s distribution relative to that of the final, largest model scale.
+In Fig.
+4c we see a sharp decrease in the W2 distance, marking the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads.
+These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal to bimodal.
 We posit that this transition marks the minimum capacity required to learn the task.
-We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply.
-We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
-2.6 Count: An Instance of U-shaped Scaling
-We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 11) but yields a very different scaling effect: a U-shaped curve.
-Figure 6 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed.
-This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 5(b).
-U-shaped scaling has been observed in LMs, but its causes are not currently well-understood.
-When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al.
-(2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
-Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution.
-Indeed, Figure 6 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width.
-The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left).
-Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
-Figure 6: Summary statistics for length generalization (count task).
-Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently.
-The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 6; however, the mean of successful runs (defined as >50% EM accuracy) 6 still improves continuously.
-(a) Qwen2.5-0.5B trained on different mixes.
-(b) Qwen2.5 models trained on 10% MMLU data.
+We also mark the point in each trend where the mode in Fig.
+3 (top left) increases sharply.
+We argue that one can draw misleading conclusions about minimal model capacity when only studying using single runs at each scale, whereas distributional statistics correctly identify the minimum capacity that a skill requires.
+3 Large Language Model Experiments
+After connecting emergence with bimodality in small synthetic settings, we turn to LMs.
+We focus on the MMLU multiple-choice question answering dataset, where performance jumps (Srivastava et al., 2023) after the LM learns the multiple choice format (Hu and Frank, 2024).
+We will show that emergent scaling curves in LMs express underlying multimodal performance distributions.
+3.1 Data and Setup
+To avoid the expense of repeatedly training large LMs from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them.
+Our continued training dataset, a mix of C4 news data and MMLU training data, is insufficient to recover the base model’s full MMLU performance.
+We hypothesize that for compute scales near the emergence threshold, performance distributions will be bimodal across random seeds.
+During pretraining, a large diverse corpus encourages models to acquire various capabilities.
+While sufficiently large models may learn all capabilities, smaller models have limited capacity, requiring capabilities to compete (Merrill et al., 2023).
+This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters.
+Task: We test LMs on MMLU (Hendrycks et al., 2021), a multiple-choice question-answering benchmark.
+Strong MMLU performance requires (1) natural language reasoning with domain knowledge and (2) producing answers in the required format.
+The latter drives emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
+Model: We use the Qwen2.5 family of base models (Yang et al., 2024). To introduce randomness, we reinitialize the final attention layer and LM head, then perform full-parameter continued pretraining.
+Data: We mix the C4 news subset (Raffel et al., 2023) with MMLU training data to ensure the multiple choice formatting circuits compete with general language modeling.
+We vary the MMLU proportion to control task-specific data size.
+Training: We continue pretraining Qwen2.5-0.5B and Qwen2.5-1.5B on C4-MMLU mixes, training 80 reinitializations per data mixture ratio.
+We train for 2 epochs (0.5B) and 5 epochs (1.5B) with learning rate 1e-5 and linear decay.
+(a) Qwen2.5-0.5B mix results.
+(b) Qwen2.5: 10% MMLU data.
 Figure 5: Performance on MMLU.
 All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU.
-Left: With insufficient MMLU training data, each model fails to follow the multiple-choice format, resulting in trivial performance.
-As more MMLU data is included, different seeds lead to bimodally distributed performance.
-With enough MMLU data, all models score above the random baseline.
-Right: For a fixed data mix, the smaller model has bimodally distributed performance.
-In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
-3 LLM Experiments
-After connecting emergence with bimodality in small synthetic settings, we turn to LMs.
-We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format (Srivastava et al., 2023; Hu and Frank, 2024).
-To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them.
-Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal.
-Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
-3.1 Data and Setup
-We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task (Srivastava et al., 2023; Snell et al., 2024).
-We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs.
-Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
-Our procedure can be regarded as continued pretraining: we mimic the model’s original pretraining objective with a dataset of English language sequences.
-We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
-During pretraining, a large diverse corpus encourages models to acquire various capabilities.
-While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023).
-This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
-Task: We test LMs on MMLU Hendrycks et al.
-(2021), a multiple-choice question-answering benchmark.
-Strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format.
-It is this latter ability which leads to emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
-Model: We use the Qwen2.5 family of base models Yang et al.
-(2024).
-To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head.
-We perform full-parameter continued pretraining on these reinitialized models.
-Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 Raffel et al.
-(2023) with the official MMLU auxiliary training data.
-The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering.
-We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
-Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 80 reinitializations on each data mixture ratio.
-We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set.
-We use a learning rate of 1e-5 with linear decay scheduler.
-We focus our analysis on experiment results from Qwen2.5-0.5B.
-In Appendix E, we include results from Qwen2.5-1.5B (Appendix 17) as well as a second task CommonSenseQA (Talmor et al., 2019) (Appendix E.2).
-3.2 Emergence Across Data Compositions
-We first examine how data composition influences the bimodal capabilities of Qwen2.5-0.5B.
-Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s performance distribution in ways that mirror the effect of scale.
-In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 Raffel et al.
-(2023), most models (out of 80 seeds) achieve near 0% performance on the MMLU test set.
-According to the existing literature, this failure stems from a failure to process the multiple-choice format Hu and Frank (2024).
-As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters.
-In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%.
-This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer.
-Finally, with sufficient data ($20\%$ MMLU), models across all 80 seeds consistently outperform the 25% random baseline.
-3.3 Emergence Across Model Scales
-Like in the synthetic setting, LM model size affects the clustered performance distribution.
-Using a training data mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 80 seeds and report MMLU test accuracy in Figure 7 (right).
-While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline.
-In fact, larger models reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales.
-This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020).
-Appendix E.2 shows similar results on the Commonsense QA (Talmor et al., 2019) task.
-Figure 8: NLL loss distributions on the MMLU test set.
-Each distribution is provided as a histogram and as a smoothed KDE.
-Even with a continuous metric (NLL), we still observe multi-modal distribution across random seeds.
-Figure 9: KDE approximation of LM training loss on 10% MMLU mixture.
-Approximation of the distribution across 80 random seeds is unimodal, smooth, and nearly symmetrical with median close to mean.
-By contrast, Figure 8 depicts MMLU loss for LMs trained on the same mix as irregularly clustered with multiple modes.
+(a) With insufficient MMLU training data, performance is trivial; as data increases, performance becomes bimodal across seeds.
+(b) For a fixed data mix, smaller models show bimodal performance while larger models consistently relearn the task, albeit below original base levels.
+3.2 Experiment Results
+Emergence Across Data Compositions.
+When training data contains more MMLU examples, the Qwen2.5-0.5B performance distribution improves in ways that mirror the effect of scale.
+In Figure 5 (left), when trained on 5% MMLU and 95% C4 (Raffel et al., 2023), most models (out of 80 seeds) achieve near 0% MMLU accuracy, failing to process the multiple-choice format.
+At 10% MMLU training data, bimodality emerges: one cluster remains near 0% accuracy (format failure), while the other reaches $\geq$25% accuracy (random baseline).
+This second cluster contains models that consistently produce valid multiple choice responses; those exceeding 25% have learned to compose format following with some recovered world knowledge.
+Due to the limited scale and quality of our continued pretraining data, fully recovering the base model’s MMLU capability is challenging.
+At $20\%$ MMLU, all 80 seeds consistently perform around or above the random baseline.
+Qwen2.5-1.5B shows the same bimodality (Figure 19), but the larger model recovers similar capabilities at much lower data mixture ratios.
+Emergence Across Model Scales.
+Model size affects the clustered MMLU performance distribution similarly to the synthetic settings in Sec.
+2.
+Using 10% MMLU training data, we continually pretrain Qwen after reinitializing with 80 different seeds (Figure 5, right).
+While Qwen2.5-0.5B forms two distinct performance clusters, Qwen2.5-1.5B consistently performs around or above the random baseline.
+Larger models reliably acquire MMLU capability on the same dataset that produces highly bimodal variation at smaller scales, consistent with scaling laws: smaller models require more training examples to match larger model performance (Rosenfeld et al., 2019; Kaplan et al., 2020).
+Figure 6: Question formation performance across data compositions.
+Left: Mean and median OOD accuracy across 80 seeds shows emergence at 10% hierarchical-rule inducing data.
+Right-4: Histograms reveal bimodal distributions at 10%, 90% and 100% hierarchical data.
+Models either cluster at either 0% (linear generalization) or 100% (hierarchical generalization).
+Figure 7: Question formation performance across model scales.
+Training on 90% hierarchical-rule inducing data, 3-layer models show unimodal distribution around 70% OOD accuracy, while 6-layer models exhibit bimodality with most seeds achieving 100% (hierarchical generalization) and some reaching 0% (linear generalization).
+Each dot represents one seed.
+Figure 8: NLL loss distributions.
+Even with a continuous metric (NLL), we still observe multi-modal distributions in performance across random seeds.
+NLL is measured on: Left: MMLU for Qwen2.5-0.5B trained on 10% MMLU mixture.
+Right: OOD QF for LM pretrained on 90% hierarchical mixture.
+Figure 9: KDE of LM training loss. The distribution across 80 seeds is unimodal and smooth. By contrast, Fig. 9 depicts MMLU loss as irregularly clustered.
 3.3 Is MMLU emergence a mirage?
 Is multiple choice performance only bimodal because accuracy is thresholded?
-As in our synthetic setting (Section 2.4), we must address the effect of the discontinuous metric.
-In Figure 8, we find that continuous negative log likelihood (NLL) loss metrics remain highly multimodal.
-The 10% MMLU ratio leads to the most even split between successes and failure in accuracy, and accordingly to clear bimodality in NLL loss, mirroring the link we observed on synthetic tasks.
-Notably, training data with a 7.5% MMLU ratio produces a performance distribution with a small “success” cluster peaked around NLL=3.4.
-The 10% MMLU ratio produces a performance distribution with a larger success cluster, also peaking around NLL=3.4.
-Therefore, the continuous metrics remain roughly bimodal—as is especially clear in the 10% mix setting, which is most evenly split between success and failure.
-While recent arguments against emergence (Schaeffer et al., 2024) claim that discontinuities disappear with continuous loss metrics, Figure 8 reveals the opposite effect in random variation—NLL loss reveals clusters that are concealed by accuracy metrics, specifically among low-performance models like those trained on the 5% MMLU mix.
-(To easily compare these metrics, see Figure 13.)
-Is this random clustering particular to the emergent task?
-As an alternative hypothesis, we might expect all tasks—including training loss—to have similarly multimodal NLL distributions, when approximated with our KDE settings.
-To the contrary, the pretraining loss produces a smooth, unimodal, and almost symmetric KDE distribution (Figure 9).
-We therefore conclude that multimodality is linked to emergent tasks—not merely a downstream effect of in-distribution training noise.
+As in our synthetic setting (Sec.
+2.4), we examine continuous metrics.
+In Fig.
+9, negative log likelihood (NLL) loss remains bimodal.
+For the 10% MMLU training data case, we confirm this bimodality is statistically significant ($p<0.001$) for significant using Hartigan’s Dip Test (Hartigan and Hartigan, 1985), which measures the maximum difference between the empirical distribution and the best-fitting unimodal distribution.
+The bimodality observed in NLL also mirrors our synthetic algorithmic task findings.
+While recent work (Srivastava et al., 2023; Schaeffer et al., 2024) argues that discontinuities disappear with continuous metrics, Fig.
+19 and Fig.
+19 reveals the exact opposite: NLL loss exposes clusters concealed by accuracy metrics, particularly among low-performance models trained on 5% MMLU (Fig.
+19) and among larger models (Fig.
+19).
+Is this clustering specific to emergent tasks?
+Fig.
+9 suggests so.
+Pretraining loss produces a smooth, unimodal, nearly symmetric distribution, contrasting sharply with MMLU’s multimodality.
+We therefore conclude that multimodality is linked to emergent tasks.
+It is not merely a downstream effect of in-distribution training noise.
+4 Grammatical Generalization in LMs
+We next examine English syntax acquisition during LM training.
+Question formation (QF) is a canonical benchmark (Mueller and Linzen, 2023; McCoy et al., 2020b, a) for grammatical rule learning.
+In this setting, the ambiguous training data supports two possible competing rules: (i) a hierarchical syntactic rule (correct for English) or (ii) a superficial linear heuristic (like an n-gram language model).
+On the disambiguating OOD test data, the hierarchical rule yields  100% accuracy, the linear rule yields  0%, and any score in between indicates inconsistent rule application.
+Prior work has documented performance clustering in QF across random seeds (Qin et al., 2024; Ahuja et al., 2024; Murty et al., 2023), with evidence that data composition and model scale influence which rule is learned (Qin et al., 2024; Ahuja et al., 2024).
+This controlled setting allows us to study bimodal distributions of competing viable solutions when training a simple language model from scratch.
+4.1 Data and Setup
+[…]
+Model: Decoder-only Transformers trained from scratch on causal language modeling.
+Following prior work (Qin et al., 2024; Murty et al., 2023), we vary depth (3-6 layers) with 8 heads per layer and 512-dimensional embeddings.
+Data: Synthetic English sentences generated via Context-Free Grammars (McCoy et al., 2018).
+Based on Qin et al.
+(2024), we vary the proportion of hierarchical-inducing versus linear-inducing sentences in the training data, paralleling our MMLU data mixture experiments.
+Training: We train 80 random seeds per configuration.
+We use Adam optimizer with learning rate $1\times 10^{-4}$ and linear decay and train models for 600K steps.
+We vary (1) model size and (2) proportion of hierarchy-inducing training examples, to study how these scale factors influence the bimodal performance distribution of hierarchical generalization.
+4.2 Experiment Results
+Emergence Across Data Compositions.
+Fig.
+6 (left) shows mean OOD accuracy exhibits sharp emergence at 10% hierarchical data, jumping from near 0% to 80%.
+However, histograms (right 4 panels) reveal that the population mean masks the underlying bimodality: at 10%, seeds split between 0% (linear rule) and 80-100% (hierarchical rule).
+Varying data composition shifts which rule models learn.
+More hierarchical-inducing sentences increase the probability of learning the hierarchical rule, but do not eliminate the competing solution even at 100% hierarchical-inducing data.
+Apparently, even our most hierarchical-inducing training set offers insufficient support to provide consistently “successful” runs, possibly because it lacks diversity.
+Emergence Across Model Scales.
+Fixing the data mixture at 90% hierarchical sentences, we now examine how model capacity affects the performance distribution.
+At 3 layers (Fig.
+7), performance is unimodal around 70%, with few seeds reaching 100%—these smaller models lack capacity to learn either systematic syntactic rule.
+At 6 layers, clear bimodality emerges: most seeds achieve 100% (hierarchical generalization) while some reach 0% (linear generalization).
+Although our largest scale models tend to consistently apply systematic rules, we find different seeds still vary in which rule they apply.
+Finally, similar to Sec.
+3.3, we confirm that the bi-modality is still observed under continuous metric in Figure 9 and is statistically significant under Hartigan’s Dip Test (Hartigan and Hartigan, 1985).
 5 Discussion
-Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities.
-The mode of our performance distributions sharply improves at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution.
-Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 2.5); the transition from the initial unimodal distribution to a bimodal one is sudden.
-Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order.
-Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018).
-Most relevant to us, the high variability of length generalization is well-documented (Zhou et al., 2024b; a).
+Our work explores the evolution of random variation in model performance across scales, bringing a nuanced perspective on emergent capabilities.
+While the mode of “emergent” performance distributions may sharply improve at a certain model scale, we attribute these sudden jumps to gradual improvements in the random distribution.
+In fact, bimodality often emerges before the mode—or most individual runs—exhibits a breakthrough, even when the transition from a unimodal distribution to a bimodal one is sudden.
+Random variation:
+Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order.
+While variation may have a benign effect on in-distribution performance (Jordan, 2024), previous works indicate that out-of-distribution performance can fluctuate substantially across training runs, even at larger scales (Madaan et al., 2024).
+Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018), including length generalization (Zhou et al., 2024b, a).
 More generally, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution loss.
@@
 (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance.
-Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation.
+Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023; Li et al., 2025) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation.
 Our work connects these random clustering effects to the phenomenon of emergence at scale.
-Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022).
-Various proposals have, consequently, aimed to improve length generalization.
-Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024).
-Others modify the dataset format by adding scratchpad or Chain-of-Thought (Anil et al., 2022) or by incorporating padding and index hints (Jelassi et al., 2023; Zhou et al., 2024a).
-Regarding random variation, Zhou et al.
+Relevant to our setting, Zhou et al.
 (2024a) and Zhou et al.
-(2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
-Emergent abilities of LMs: In LLMs, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020).
-These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023).
-Understanding the conditions and mechanisms underlying emergence is a key area of research.
-Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Srivastava et al., 2023; Schaeffer et al., 2024), some breakthrough capabilities remain emergent.
+(2024b) provided evidence of variability in length generalization across random seeds, which we further analyze across model scales.
+We also expand on the findings of Qin et al.
+(2024) of bimodal variation in LM syntax acquisition.
+Emergent abilities of LMs:
+In LMs, scaling laws predict reliable performance improvements as models increase in size or train on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020).
+Emergent abilities are abilities that arise unexpectedly, out of line with these predictions (Srivastava et al., 2023; Ganguli et al., 2022; Wei et al., 2022).
+These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales.
+Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior (Srivastava et al., 2023; Schaeffer et al., 2024), other breakthrough capabilities remain emergent—8% of high-breakthroughness datasets in BIG-Bench, according to Schaeffer et al.
+(2024).
+Our findings suggest that careful statistical analysis is required to evaluate claims both for and against “emergence” scaling on specific tasks.
 Snell et al.
@@
 Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
-Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022).
-For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024).
-Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al.
-(2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently.
+Depth versus Width Scaling:
+Downstream performance varies with architecture shape, not just model size (Tay et al., 2022).
+For compositional tasks, deeper models often generalize better, but for fixed compute budgets, shallower and wider models may be advantageous (Petty et al., 2024).
+Several explanations of this benefit have been proposed.
+Edelman et al.
+(2024) argued that wider networks offer more parallel queries over randomized subnetworks and therefore learn sparse features efficiently.
 Levine et al.
-(2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns.
-In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
-In one surprising result, we document a regime in which increasing width damages model performance but increasing depth improves model performance.
-In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
+(2020) established width-dependent depth thresholds beyond which depth yields diminishing returns (Levine et al., 2020).
+We investigate how independently scaling width and depth influences random variation in compositional tasks, documenting a surprising regime (Appx.
+C.2) where increasing width damages performance while increasing depth improves it.
+Our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
 Acknowledgments
@@
 Our work has been improved by invaluable discussion with Will Merrill and David Chiang.
-Reproducibility Statement
-Upon publication, we will release the code for both synthetic and LLM experiments, including code, and data.
-Detailed descriptions of model, data and training are provided in Appendix A, and Section 3.
 Appendix A Experimental details on synthetic tasks
-Below we provide more details when training decoder-only language models on the count and addition—with index hints—tasks.
-Hyperparameters are largely adapted from [Zhou et al., 2024a].
+Below we provide more details on the count and addition task settings.
+Hyperparameters for our decoder-only Transformer models are largely adapted from (Zhou et al., 2024a).
 We train all of our models to convergence on the train distribution.
 Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
-Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
-  • Model scales: As mentioned in Section 2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width.
+At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
+Count task: For all of our training runs, we fix the vocabulary size to 150.
+For thresholded evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
+  • Model scales: As mentioned in Sec.
+2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width.
 The precise parameters for each variation are as follows.
@@
 We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
-Reverse Order Addition with Index Hints: For evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.
+Reverse Order Addition with Index Hints: For thresholded evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.
   • Model scales: For our fixed depth experiments, we fix the network depth to 6 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768\}$.
@@
 Srivastava et al.
-[2023] introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale.
+(2023) introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale.
 Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
-$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})},$
-$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$
+$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})}$ (2)
+$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$ (3)
 where $I(y)=\mathrm{sign}(\arg\max_{i}y_{i}-\arg\min_{i}y_{i})(\max_{i}y_{i}-\min_{i}y_{i})$.
-In Figure 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and addition (below).
+In Fig. 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for (a) count and (b) addition.
 (a) Count.
 (b) Addition.
-Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 10(a) count and 10(b) addition.
-Appendix C Additional Exact Match Accuracy Histogram Plots
+Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appx. B, with resulting seeds plotted for 10(a) count and 10(b) addition.
+Appendix C Count task
+Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number.
+Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training.
+Zhou et al.
+(2024a) showed that models trained to count can generalize to more than twice this training length; however, Appx.
+C.2 reveals a more nuanced view of length generalization based on its distribution across independent model runs.
+C.1 Count performance distributions
 (a) Fixing network depth to 4 layers.
 (b) Fixing network width to 4 heads per layer.
-Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling 11(a) width and 11(b) depth.
+Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling (11(a)). width and (11(b)). depth.
 Figure 12: Histograms for EM accuracy when fixing depth to be one layer (count task).
@@
 While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
-Our plots in the main paper fixed test length 40 for addition (Figure 2).
-In Figure 11 we show analogous histograms for a fixed test length of 60 on the count task.
+Our plots in the main paper show distributions of addition task accuracy with a fixed test length 40 (Fig.
+2).
+In Fig.
+11 we show analogous histograms for the count task with a fixed test length of 60.
 We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
-In Section 2.5, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task.
-For example, Figure 12 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
+In Sec.
+2.5, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task.
+For example, Fig.
+12 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
+C.2 Count task as a case of U-shaped random scaling
+We next consider the counting task (Sec.
+2.1), which also exhibits bimodally-distributed performance (see Appx.
+Fig.
+11) but yields a very different scaling effect: a U-shaped curve.
+Fig.
+14 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed.
+This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Fig.
+13.
+U-shaped scaling has been observed in LMs, but its causes are not currently well-understood.
+When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al.
+(2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models.
+Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution.
+Indeed, Fig.
+14 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width.
+The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left).
+Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
+Figure 13: Changes in random variation (count task).
+Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently.
+We mark the emergence of bimodality at the last scale before multiple peaks appear.
+We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
+The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Fig.
+14.
+Figure 14: As in Fig.
+3, we capture summary statistics for test length 60 on the count task.
+For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case.
+The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
 Appendix D Additional Continuous Metric Plots
-In Section 2.4 we have seen seen that the addition histograms are bimodal in exact match accuracy (Figure 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Figure 4).
+Addition Task.
+In Sec.
+2.4 we have seen seen that the addition histograms are bimodal in exact match accuracy (Fig.
+2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Fig.
+4 a, b ).
 We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
@@
 Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric.
-We plot the resulting histograms of this metric in Figures 14 and 15 for addition and count respectively.
+We plot the resulting histograms of this metric in Fig.s 15 and 16 for addition and count respectively.
 We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
-In Figure 16, we plot the analogous KDE plot of the loss-based error metric for the count task.
+Count Task.
+In Fig.
+17, we plot the analogous KDE plot of the loss-based error metric for the count task.
 For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition.
-However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 11 and 15 is still associated with stronger bimodality in the loss plots.
+However, we claim that stronger bimodality in the original accuracy/probability plots in Fig.s 11 and 16 is still associated with stronger bimodality in the loss plots.
 For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions.
 For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
-Finally, Figure 13 provides a copy of Figure 8 with both accuracy and NLL loss comparable for easy contrast.
+Large LM Experiments.
+Finally, Fig.
+19 provides a copy of Fig.
+9 with both accuracy and NLL loss comparable for easy contrast.
 We note that high-performance clusters are distinguishable in both accuracy and loss, but the thresholded accuracy conceals low-performance clusters which are visible in the continuous distribution.
-Figure 13: KDE approximations of performance variation on the MMLU test set.
-Provided for contrasting accuracy and continuous NLL loss.
-Top: Approximate distribution of accuracy metrics.
-Bottom: Approximate distribution of KDE metrics, copied from Figure 8.
+We also include histogram version of results in Fig.
+19.
+Due to computation limit, for each data mixes, we train on 20 random seeds.
+Similar to Fig.
+19.
+models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
 (a) Fixing network depth to 6 layers while scaling up width.
@@
 Metrics are calculated by Equation 4.
-The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Figure 4.
+The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Fig.
+4 (a, b).
 (a) Fixing network depth to 4 layers.
@@
 Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs.
-At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
-Appendix E Additional Language Model Results
-E.1 MMLU Results on Qwen2.5-1.5B
-We repeat the same MMLU experiment (Section 3.1) on Qwen2.5-1.5B and we include results in Figure 17.
-Due to computation limit, for each data mixes, we train on 20 random seeds.
-Similar to Figure 13.
-models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
+At scales where the EM accuracy distribution is most strongly bimodal (width=4, width=8, and depth=2), KDE places the most density at areas of high loss.
+Figure 18: KDE approximations of performance variation on the MMLU test set.
+Provided for contrasting accuracy and continuous NLL loss.
+Top: Approximate distribution of accuracy metrics.
+Bottom: Approximate distribution of KDE metrics, copied from Fig.
+9.
 Figure 19: Accuracy and NLL loss for Qwen2.5-1.5B MMLU experiments.
@@
 Bottom: Distribution of NLL loss metric.
-E.2 CSQA LM Results
-In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA, Talmor et al.
-[2019]) task and we include the result in Figure 18.
-Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
-Figure 18: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.
+Appendix E Grammatical Generalization Details
+E.1 Task Description
+In the question formation (QF) task, models transform declarative English sentences into questions by moving an auxiliary verb to the front of the sentence.
+The task tests whether models learn the correct hierarchical syntactic rule based on the sentence’s syntax tree structure, or instead rely on a superficial linear heuristic.
+Our training data (based on McCoy et al. (2018)) permits two strategies for choosing which verb to move:
+  1. 1.
+    Linear rule: Move the first auxiliary verb in the sentence (incorrect for English)
+  2. 2.
+    Hierarchical rule: Move the main auxiliary verb, determined by the sentence’s syntactic structure (correct for English)
+Examples of each rule are provided in Table 1.
+The first example is ambiguous because both the hierarchical and linear rules produce the same correct output.
+In contrast, the second example is unambiguous because only the hierarchical rule produces the grammatically correct question.
+E.2 Data Composition
+The training data contains two components:
+  1. 1.
+    Question formation examples: Transform declarative sentences to questions. These examples are always ambiguous between linear and hierarchical rules.
+  2. 2.
+    Declaration copying examples: Simply repeat declarative sentences without transformation. Unlike QF examples, these can include unambiguous sentence structures.
+Following Qin et al.
+(2024), we manipulate the declaration copying subset to control exposure to different syntactic structures.
+Specifically, we vary the ratio of center-embedded sentences versus right-branching sentences:
+  • Center-embedded sentences contain recursive clause embedding (e.g., “My unicorn who doesn’t sing does move”), which requires hierarchical syntactic processing and encourages learning the hierarchical rule.
+  • Right-branching sentences have simpler linear structure (e.g., “My unicorn does move the dogs that do wait”), which can be solved with linear n-gram patterns.
+We create training sets with different mixtures of these sentence types in the declaration copying subset: Quest Only (no declaration copying), Center Embed (only center-embedded declarations), Right Branch (only right-branching declarations), and various intermediate mixtures.
+E.3 Evaluation
+The in-distribution test set contains only ambiguous examples (solvable by both rules), while the OOD test set contains only unambiguous examples (requiring the hierarchical rule).
+Models using the hierarchical rule achieve $\sim$100% accuracy on both test sets.
+Models using the linear rule achieve $\sim$100% on in-distribution data but $\sim$0% on OOD data.
+We therefore use OOD accuracy as our measure of whether models have acquired the hierarchical syntactic rule.
+Table 1: Question Formation task examples.
+Models must move the main auxiliary verb to the front to form a question.
+Ambiguous examples can be solved by either linear or hierarchical rules; unambiguous examples require the hierarchical rule.
+Task Example Type Examples
+Question Formation Ambiguous Input: My unicorn does move the dogs that do wait.
+Output: Does my unicorn move the dogs that do wait?
+Unambiguous Input: My unicorn who doesn’t sing does move.
+Linear Output: Doesn’t my unicorn who sing does move?
+Hierarchical Output: Does my unicorn who doesn’t sing move?